### v2

# An Optimal Design Framework for Lasso Sign Recovery

Abstract
Supersaturated designs investigate more factors than there are runs, and are often constructed under a criterion measuring a design’s proximity to an unattainable orthogonal design. The most popular analysis identifies active factors by inspecting the solution path of a penalized estimator, such as the lasso. Recent criteria encouraging positive correlations between factors have been shown to produce designs with more definitive solution paths so long as the active factors have positive effects. Two open problems affecting the understanding and practicality of supersaturated designs are: (1) do optimal designs under existing criteria maximize support recovery probability across an estimator’s solution path, and (2) why do designs with positively correlated columns produce more definitive solution paths when the active factors have positive sign effects? To answer these questions, we develop criteria maximizing the lasso’s sign recovery probability. We prove that an orthogonal design is an ideal structure when the signs of the active factors are unknown, and a design constant small, positive correlations is ideal when the signs are assumed known. A computationally-efficient design search algorithm is proposed that first filters through optimal designs under new heuristic criteria to select the one that maximizes the lasso sign recovery probability.
Keywords: Screening experiments; constrained-positive $Var(s)$-criterion; variable selection; $UE(s^{2})$-criterion; Gauss Dantzig selector
1 Introduction
A screening experiment aims to learn which $k$ out of $p$ factors most influence the response variable using as few runs, $n$, as possible. Achieving this goal with a traditional least-squares analysis is challenging because often not all possible factorial effects can estimated simultaneously. Still, effective screening is feasible under the effect sparsity principle and when the experimental design is carefully selected (Box and Hunter, 1961a, b; Mee, 2009; Xu et al., 2009; Mee et al., 2017). Two-level supersaturated designs, or SSDs, push screening to its limits with $n<p+1$ so that not even the $p$ main effects can be estimated simultaneously via least-squares. The recent literature on SSD analysis instead promotes penalized regression methods to perform screening under the main effects model, but the literature on the construction of SSDs has not kept up with this development. The present work addresses the gap between analysis and construction by developing an optimal design framework that tailors design selection to statistical properties of the lasso estimator.
Throughout this paper, we will assume a main effects model $\boldsymbol{y}=\beta_{0}\boldsymbol{1}+\textbf{X}\boldsymbol{\beta}+\boldsymbol{e}$ where X is the $n\times p$ design matrix with elements $x_{ij}=\pm 1$, $\boldsymbol{\beta}=(\beta_{1},\dots,\beta_{p})^{T}$ is a sparse vector with $k<p$ nonzero elements, and $\boldsymbol{e}\sim N(\boldsymbol{0},\sigma^{2}\textbf{I})$. Without loss of generality, assume $\sigma^{2}=1$, making $\boldsymbol{\beta}$ the signed signal-to-noise ratios. Then the analysis goal is estimation of the support of $\boldsymbol{\beta}$, denoted $\mathcal{A}=\{j:|\beta_{j}|>0\}$, although for technical reasons we will be more interested in estimating the sign vector of $\boldsymbol{\beta}$, denoted $\boldsymbol{z}=\text{sign}(\beta_{j})$. We will focus our attention on optimal SSDs, although the framework developed in this paper could be applied to more general screening designs. Note that when we indicate designs are “optimal”, we are referring to the best designs found by heuristic optimality procedures, unless otherwise indicated.
1.1 Background and Open Problems
The least-squares estimator for $\boldsymbol{\beta}$ is not unique for any SSD, complicating its analysis and design construction. If there were SSDs with unique least-squares estimators, the ideal X would satisfy $\textbf{S}=\textbf{L}^{T}\textbf{L}=n\textbf{I}_{p+1}$ where $\textbf{L}=(\boldsymbol{1}|\textbf{X})$, as its $\hat{\beta}_{j}$ would have the minimum possible variance across all SSDs. These so-called orthogonal designs only exist when $n=0\,(\text{mod}\ 4)>p$. For arbitrary $n>p$, a design could instead be selected by minimizing a variance-based criterion such as the $D$-criterion, $|(\textbf{L}^{T}\textbf{L})^{-1}|$, or $A$-criterion, $\text{tr}[(\textbf{L}^{T}\textbf{L})^{-1}]$. This optimal design framework based on least-squares is well-developed and tractable (Pukelsheim, 2006; Goos and Jones, 2011), but is not applicable for SSDs.
Most SSD construction methods focus on optimizing heuristic criteria that find a design whose S is close to $n\textbf{I}_{p+1}$. Booth and Cox (1962) proposed the criterion $E(s^{2})=\sum_{1<i<j\leq p}s_{ij}^{2}$ assuming balanced designs (i.e., $\textbf{X}^{T}\boldsymbol{1}=\boldsymbol{0}$), which has been extensively studied (see Lin, 1993; Wu, 1993; Nguyen, 1996; Li and Wu, 1997; Tang and Wu, 1997; Ryan and Bulutoglu, 2007). Jones and Majumdar (2014) proposed the unconditional $E(s^{2})$-criterion $UE(s^{2})=\sum_{0\leq i<j\leq p}s_{ij}^{2}$ that includes the $s_{0j}^{2}$ elements. Other criteria have been proposed that measure a design’s quality in terms of subsets of columns (Deng et al., 1996, 1999; Jones et al., 2009), which is connected to a stepwise or all-subsets analysis (Lin, 1995; Abraham et al., 1999; Westfall et al., 1998; Liu et al., 2007). Chen and Lin (1998) and Sarkar et al. (2009) investigated support recovery properties of SSDs under such sequential analyses.
Li and Lin (2002) and Phoa et al. (2009) advocated for SSDs to be analyzed under penalized estimation that induces sparse estimates. The previously mentioned SSD criteria, however, are not directly tied to statistical properties of such an analysis. There have been some attempts in the literature to address this disconnect. Marley and Woods (2010), Draguljić et al. (2014), and Weese et al. (2015) have compared optimal SSDs under different criteria via simulation with the lasso and/or Gauss-Dantzig selector (GDS), finding little difference between the SSDs in terms of support recovery. Recently, Weese et al. (2017) and Weese et al. (2021) proposed $Var(s+)$-optimal designs that minimize
$Var(s)=UE(s^{2})-UE(s)^{2}\text{ such that }\frac{UE^{*}(s^{2})}{UE(s^{2})}>c\text{ and }UE(s)>0\ ,$
where $UE^{*}(s^{2})$ is the optimal $UE(s^{2})$ value for the given $n$ and $p$, and $UE(s)=\sum_{0\leq i<j\leq p}s_{ij}$. The ideal structure of S under $Var(s+)$ is notably non-orthogonal, having average off-diagonal elements equal to a small, positive constant, making $Var(s)=0$ while satisfying the constraints on $UE(s^{2})$ and $UE(s)$. Simulation studies showed $Var(s+)$-optimal designs (with $c=0.8$) performed similarly to other optimal SSDs under the GDS, but when the effect directions were known in advance (i.e. the nonzero signs of $\boldsymbol{\beta}$ were all positive), the $Var(s+)$-optimal design had better support recovery properties. The simulation studies mentioned above only perform relative comparisons of SSDs generated by heuristic criteria and so do not answer the following open problems:
  1. 1.
    How well do existing heuristic criteria find SSDs that maximize the probability of support recovery across all possible SSDs?
  2. 2.
    Why does the $Var(s+)$-criterion identify SSDs with better support recovery probability when the active effects have positive signs?
The theory developed in this paper answers both of these problems.
Singh and Stufken (2022) addressed problem 1 by proposing two criteria related to the restricted isometry property (Candes and Tao, 2007) and irrepresentable condition (Gai et al., 2013) of the GDS, which are sufficient for estimation and sign recovery consistency. For given $\mathcal{A}$ of size $k=|\mathcal{A}|$, let $\textbf{C}_{\mathcal{A}}$ denote the correlation matrix of the columns in $\mathcal{A}$. Let $0\leq\lambda_{\mathcal{A}1}\leq 1$ and $\lambda_{\mathcal{A}k}\geq 1$ denote the minimum and maximum eigenvalues of $\textbf{C}_{\mathcal{A}}$, respectively, and define $\tau_{\mathcal{A}}=\max(1-\lambda_{\mathcal{A}1},\lambda_{\mathcal{A}k}-1)$. Singh and Stufken (2022) propose calculating the 95th percentile of all $\tau_{\mathcal{A}}$, denoted by $\zeta_{k}$. They then calculate $\zeta_{k}$ for $k=2,\dots,\lceil n/2\rceil$ and define the estimation capacity criterion:
$\text{EC}=\text{mean}(\zeta_{2},\dots,\zeta_{\lceil n/2\rceil})\ .$
Ideally, EC would be 0, corresponding to an orthogonal design. Their second criterion assumes the sign of the active effects is always positive, denoted by $\boldsymbol{1}_{k}$. First, for a given $\mathcal{A}$, the following two weak sign consistency conditions are checked
$|\textbf{C}_{\mathcal{I}\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}_{k}|$ $<\boldsymbol{1}_{p-k}$ (1)
$|\textbf{C}_{\mathcal{I}\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\text{sign}(\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}_{k})|$ $<\boldsymbol{1}_{p-k}\ ,$ (2)
where $\textbf{C}_{\mathcal{I}\mathcal{A}}$ is the matrix of cross correlations of columns between the inactive set, =$\mathcal{I}=\{j:\beta_{j}=0\}$, and $\mathcal{A}$. Similar to EC, this condition is checked across all $\mathcal{A}$ where $|\mathcal{A}|=k$ and $\gamma_{k}$ denotes the proportion of such $\mathcal{A}$ that achieve both conditions. Finally, their weak sign consistency criterion is defined to be
$\text{WSC}=\text{mean}(\gamma_{2},\dots,\gamma_{\lceil n/2\rceil})\ .$ (3)
We want to find a design that maximizes WSC. It would be preferred that the left hand sides of (1) and (2) equal 0, again corresponding to orthogonal designs, but nonorthogonal designs may still have a large WSC. The authors employ a hybrid elitist Pareto optimality-based coordinate exchange algorithm (Cao et al., 2017) to identify designs on the Pareto front of the EC and WSC. We refer to their optimal designs as Dantzig-based Consistency Designs (DCDs).
The DCD criteria require evaluations across all $\mathcal{A}$ of sizes $k=2,\dots,\lceil n/2\rceil$. To ease the computational burden, Singh and Stufken (2022) randomly sample subsets for each $k$. Furthermore, their examples initialized their search algorithms with only two designs each, a $Var(s+)$-optimal design and either a random, $E(s^{2})$-optimal, or $UE(s^{2})$-optimal design. The resulting DCDs initialized at a $Var(s+)$ design were shown via simulation to have better support recovery properties than the $Var(s+)$-optimal design assuming positive signs of the active effects. However, their improvement over the $Var(s+)$-optimal design may be attributed to their initialization at that design. That is, they show that the initialized $Var(s+)$-optimal design is not globally optimal for known signs, but there is no guarantee a DCD is either. Indeed, an orthogonal design is the ideal design under their criteria. It may be that the globally optimal DCD performs worse than the $Var(s+)$ design. Still, their results provide strong support for constructing SSDs under criteria directly tied to statistical properties of penalized estimators.
Others have considered the role of X in the statistical properties of the lasso, though not in the context of SSDs. Zhao and Yu (2006) identified the strong irrepresentable condition (SIC) on X for establishing support recovery consistency of the lasso (Zhang and Huang, 2008; Jia and Rohe, 2015). Wainwright (2009) and Hastie et al. (2019) have studied support recovery for random designs where the rows of X are independently generated from a zero-mean multivariate Normal distribution. Random designs, however, generally perform poorly for small run designs like SSDs. Based on the lasso’s SIC, Deng et al. (2013) constructed SSDs from nearly orthogonal Latin hypercube designs to minimize the off-diagonals of S assuming factors have settings between $[0,1]$. This is essentially a construction technique for minimizing $UE(s^{2})$ with more general factor settings, even though SSDs commonly assume fixed settings of ±$\pm 1$. Xing (2015) proposed a lasso SIC criterion to construct optimal two-level SSDs, similar to the $\gamma_{k}$ from (3). Huang et al. (2020) proposed a lasso optimal design theory that applies variance-based criteria to the approximate covariance matrix of the debiased lasso (Javanmard and Montanari, 2014), which is capable of performing inference via confidence intervals. Under the framework of approximate designs (i.e., $n\to\infty$) they note their criteria are not convex and give an equivalence theorem for establishing whether a design is locally optimal. They then propose an algorithmic construction for generating many local optimal approximate designs and implement a rounding procedure on the approximate design’s replication weights to produce an exact design. Overall, their approach requires many stages of approximation, leading to a discrepancy between the approximate and exact design’s covariance matrix.
1.2 Contributions and Overview
This paper addresses the two SSD open problems by developing optimality criteria that directly maximize the probability of sign recovery for the lasso estimator. The Dantzig selector and its statistical properties are closely related to the lasso estimator (Meinshausen et al., 2007; Lounici, 2008; Bickel et al., 2009; James et al., 2009; Asif and Romberg, 2010). The statistical properties of the lasso are tractable, making its optimal design framework more straightforward.
The first major contribution of this paper is a local optimal design approach to maximize the sign recovery probability assuming $\boldsymbol{\beta}$ is known and the lasso tuning parameter, $\lambda>0$, is fixed. We then develop criteria that relax the model assumptions and summarize the sign recovery probability across a range of $\lambda$. As optimization of the proposed criteria is challenging, our second major contribution is a criterion based on the $p\times p$ correlation matrix of X. We show the ideal correlation matrices under different heuristic criteria are optimal under our new criteria, thus providing solutions to both open problems. In particular, we prove the surprising result that an orthogonal correlation matrix is suboptimal when $\boldsymbol{z}$ is known. Our final contribution is a computationally efficient construction algorithm that leverages the speed of exchange algorithms for optimizing heuristic criteria.
The paper is organized as follows. Section 2 develops exact design optimality criteria targeting the lasso’s sign recovery probability under different assumptions about the model. Section 3 presents a pseudo-approximate optimal design framework that targets the optimal correlation matrix of X and justifies the ideal designs sought after by heuristic criteria under different assumptions about $\boldsymbol{z}$. Section 4 describes a computationally efficient algorithm for constructing exact optimal designs under our proposed criteria. We then demonstrate the optimal SSDs under our new framework in Section 5. We conclude the paper with a discussion in Section 6, describing important implications of our results and future work.
2 Exact Local Optimality Criteria
The lasso is often applied to the centered and scaled design matrix $\textbf{F}=(\textbf{I}-\textbf{P}_{1})\textbf{X}\textbf{V}^{-1/2}$ where $\textbf{P}_{1}=n^{-1}\boldsymbol{1}\boldsymbol{1}^{T}$ and V is a diagonal matrix comprised of the diagonal elements of $n^{-1}\textbf{X}^{T}(\textbf{I}-\textbf{P}_{1})\textbf{X}$. The analysis then targets support/sign recovery of $\boldsymbol{\beta}^{*}=\textbf{V}^{1/2}\boldsymbol{\beta}$. The diagonal elements of $\textbf{V}^{1/2}$ are nonnegative and bounded above by 1, making $|\beta^{*}_{j}|\leq|\beta_{j}|$, and $\text{sign}(\beta_{j}^{*})=\text{sign}(\beta_{j})$ when V has all positive diagonal elements. The support, $\mathcal{A}$, is estimated by the support of the lasso estimator
$\hat{\boldsymbol{\beta}}^{*}=\operatorname*{arg\,min}_{\boldsymbol{\beta}^{*}}\frac{1}{2}\boldsymbol{\beta}^{*T}\textbf{C}\boldsymbol{\beta}^{*}-\frac{1}{n}\boldsymbol{y}^{T}\textbf{F}\boldsymbol{\beta}^{*}+\lambda\sum_{j=1}^{p}|\beta_{j}^{*}|\ ,$
where $\textbf{C}=n^{-1}\textbf{F}^{T}\textbf{F}$ is the correlation matrix of the columns of X. Denote the estimated support and sign of the resulting lasso estimate, $\hat{\boldsymbol{\beta}}$, by $\hat{\mathcal{A}}=\{j:|\hat{\beta}_{j}|>0\}$ and $\hat{\boldsymbol{z}}=\text{sign}(\hat{\beta}_{j})$, respectively.
We denote submatrices of X and F corresponding to column subsets $\mathcal{T}\subseteq\{1,\dots,p\}$ by $\textbf{X}_{\mathcal{T}}$ and $\textbf{F}_{\mathcal{T}}$, respectively. For a $p\times 1$ vector, $\boldsymbol{v}$, $\boldsymbol{v}_{\mathcal{T}}$ denotes the $|\mathcal{T}|\times 1$ vector with the $\mathcal{T}$ elements of $\boldsymbol{v}$. For all other matrices, we will consider submatrices by selecting both rows and columns with two index sets $\mathcal{U}$ and $\mathcal{T}$. That is, for a matrix M, let $\textbf{M}_{\mathcal{U}\mathcal{T}}$ denote the submatrix of M with rows and columns indexed by $\mathcal{U}$ and $\mathcal{T}$, respectively. For brevity, we will denote $\textbf{M}_{\mathcal{T}\mathcal{T}}=\textbf{M}_{\mathcal{T}}$, which should not be confused with a subsetting of columns alone.
For known $\hat{\mathcal{A}}$ and $\hat{\boldsymbol{z}}$, the lasso KKT conditions are (Zhao and Yu, 2006; Tibshirani, 2012):
$\hat{\boldsymbol{\beta}}_{\hat{\mathcal{A}}}^{*}$ $=\frac{1}{n}\textbf{C}_{\hat{\mathcal{A}}}^{-1}\textbf{F}_{\hat{\mathcal{A}}}^{T}\boldsymbol{y}-\lambda\textbf{C}_{\hat{\mathcal{A}}}^{-1}\hat{\boldsymbol{z}}_{\hat{\mathcal{A}}}\ ,$ (4)
$\boldsymbol{0}$ $<\hat{\textbf{Z}}_{\hat{\mathcal{A}}}\hat{\boldsymbol{\beta}}_{\hat{\mathcal{A}}}^{*}\ ,$ (5)
$\ \lambda\boldsymbol{1}$ $\geq\left|\textbf{C}_{\hat{\mathcal{I}}\hat{\mathcal{A}}}\hat{\boldsymbol{\beta}}_{\hat{\mathcal{A}}}^{*}-\frac{1}{n}\textbf{F}_{\hat{\mathcal{I}}}^{T}\boldsymbol{y}\right|\ ,$ (6)
where $\textbf{C}_{\hat{\mathcal{A}}}^{-1}=n(\textbf{F}_{\hat{\mathcal{A}}}^{T}\textbf{F}_{\hat{\mathcal{A}}})^{-1}$, $\hat{\textbf{Z}}_{\hat{\mathcal{A}}}=\text{Diag}(\hat{\boldsymbol{z}}_{\hat{\mathcal{A}}})$, and ^=$\hat{\mathcal{I}}=\{j:\hat{\beta}_{j}=0\}$. For given X, $\boldsymbol{y}$, and $\boldsymbol{\beta}$, we can check whether the resulting $\hat{\boldsymbol{\beta}}^{*}$ has the true support and sign vector by setting $\hat{\mathcal{A}}=\mathcal{A}$ and $\hat{\boldsymbol{z}}_{\hat{\mathcal{A}}}=\boldsymbol{z}_{\mathcal{A}}$ and checking the KKT conditions. If they hold, $\hat{\boldsymbol{\beta}}^{*}$ recovers the sign, which is more stringent than recovering the support. The proposed criteria in this section rank designs according to the probability of sign recovery: $P(\hat{\boldsymbol{z}}=\boldsymbol{z}\,|\,\textbf{F},\,\boldsymbol{\beta})$, which depends on the unknown parameter values and $\lambda$. Therefore, to calculate the probability, one must specify a $\lambda>0$ and a $\boldsymbol{\beta}$, so our proposed framework falls in the class of local optimal designs commonly employed with nonlinear models and maximum likelihood estimation (Silvey, 1980; Khuri et al., 2006; Yang and Stufken, 2009). This should not be particularly surprising because the probability is similar to a power calculation, which requires specification of effect sizes. We will recall this connection in Section 2.2 where we introduce more practical criteria.
Clearly $P(\hat{\boldsymbol{z}}=\boldsymbol{z}\,|\,\textbf{F},\,\boldsymbol{\beta})$ is a joint probability of two events. The first event follows from equations (4) and (5) and checks whether $\hat{\boldsymbol{\beta}}_{\mathcal{A}}$ has the sign $\boldsymbol{z}_{\mathcal{A}}$:
$S_{\lambda}\,|\,\textbf{F}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}}$ =$=\left\{\boldsymbol{u}<\sqrt{n}\textbf{Z}_{\mathcal{A}}\textbf{V}^{1/2}_{\mathcal{A}}\boldsymbol{\beta}_{\mathcal{A}}\right\}\ ,$ (7)
where $\boldsymbol{u}\sim N(\lambda\sqrt{n}\,\textbf{Z}_{\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{z}_{\mathcal{A}},\textbf{Z}_{\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\textbf{Z}_{\mathcal{A}})$. Note we define the event with respect to $\boldsymbol{\beta}_{\mathcal{A}}$ rather than the design-dependent $\boldsymbol{\beta}_{\mathcal{A}}^{*}$. The second event follows from equations (4) and (6) and coincides with all $j\in\mathcal{I}$ having $\hat{\beta}^{*}_{j}=0$:
$I_{\lambda}\,|\,\textbf{F},\,\boldsymbol{z}=\left\{|\boldsymbol{v}|\leq\lambda\sqrt{n}\boldsymbol{1}\right\}\ ,$ (8)
where $\boldsymbol{v}~{}N(\lambda\sqrt{n}\textbf{C}_{\mathcal{I}\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{z}_{\mathcal{A}},\textbf{C}_{\mathcal{I}}-\textbf{C}_{\mathcal{I}\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\textbf{C}_{\mathcal{A}\mathcal{I}})$. The event depends on $\boldsymbol{\beta}$ only though $\boldsymbol{z}$.
Remark 1 There is a tradeoff between the probabilities of (7) and (8) so both should be considered simultaneously. Deng et al. (2013) and Xing (2015) focus on heuristic criteria based on the lasso’s SIC, which optimizes (8) alone.
Remark 2 The probability of support recovery, $P(\hat{\mathcal{A}}=\mathcal{A}\,|\,\textbf{F},\,\boldsymbol{\beta})$, requires consideration of all possible $2^{k}$ sign vectors that have the same 00 elements as $\boldsymbol{z}$ but alternative ±$\pm 1$ elements indexed by $\mathcal{A}$. Defining $\mathcal{Z}_{\mathcal{A}}$ as the set of all such sign vectors, we have $P(\hat{\mathcal{A}}=\mathcal{A}\,|\,\textbf{F},\,\boldsymbol{\beta})=\sum_{\tilde{\boldsymbol{z}}\in\mathcal{Z}_{\mathcal{A}}}P(\hat{\boldsymbol{z}}=\tilde{\boldsymbol{z}}\,|\,\textbf{F},\,\boldsymbol{\beta})$. Each $P(\hat{\boldsymbol{z}}=\tilde{\boldsymbol{z}}\,|\,\textbf{F},\,\boldsymbol{\beta})$ may be calculated by replacing $\boldsymbol{z}_{\mathcal{A}}$ and $\boldsymbol{z}$ with $\tilde{\boldsymbol{z}}_{\mathcal{A}}$ and $\tilde{\boldsymbol{z}}$, respectively, in events (7) and (8).
Remark 3 As $|\boldsymbol{\beta}_{j}^{*}|\to\infty$ for all $j\in\mathcal{A}$, the probability of sign recovery and support recovery are equivalent. To see this, note that $\textbf{Z}_{\mathcal{A}}\textbf{V}^{1/2}_{\mathcal{A}}\boldsymbol{\beta}_{\mathcal{A}}=\textbf{V}^{1/2}_{\mathcal{A}}|\boldsymbol{\beta}_{\mathcal{A}}|$ so $P(S_{\lambda}\,|\,\textbf{F}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})\to 1$. For any other $\tilde{\boldsymbol{z}}\in\mathcal{Z}_{\mathcal{A}}$, $\tilde{\textbf{Z}}_{\mathcal{A}}\textbf{V}^{1/2}_{\mathcal{A}}\boldsymbol{\beta}_{\mathcal{A}}$ would have negative elements, causing $P(S_{\lambda}\,|\,\textbf{F}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})\to 0$. Finally, as $I_{\lambda}$ is independent of $|\boldsymbol{\beta}_{\mathcal{A}}|$, $P(\hat{\mathcal{A}}=\mathcal{A}\,|\,\textbf{F},\,\boldsymbol{\beta})\to P(\hat{\boldsymbol{z}}=\boldsymbol{z}\,|\,\textbf{F},\,\boldsymbol{\beta})$.
Remark 4 We assume $\textbf{C}_{\mathcal{A}}^{-1}$ always exists, making $\boldsymbol{u}$ nondegenerate, but $\boldsymbol{v}$ will have a degenerate multivariate Normal distribution for SSDs because C cannot be full rank. The probability of this event can be calculated following some linear transformation of $\boldsymbol{v}$.
2.1 Local criteria assuming known $\boldsymbol{\beta}$
For a fixed $\lambda$ and known $\boldsymbol{\beta}$, define the local optimality criterion
$\phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta})=P(\hat{\boldsymbol{z}}=\boldsymbol{z}\,|\,\textbf{F},\,\boldsymbol{\beta})=P(S_{\lambda}\cap I_{\lambda}\,|\,\textbf{F},\,\boldsymbol{\beta})\ .$ (9)
A $\phi_{\lambda}$-optimal design is $\textbf{X}^{*}=\text{argmax}_{\textbf{X}}\ \phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta})$. This approach is impractical, particularly due to its perfect knowledge of $\mathcal{A}$, but it is foundational for the more practical criteria that allow for uncertainty about $\mathcal{A}$. The following is a fundamental result about the role of $\boldsymbol{z}$ in $\phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta})$. Its proof and all future proofs may be found in the Supplementary Materials.
Lemma 1.
For a given X and its F, consider a $\boldsymbol{\beta}$ and its reflection $-\boldsymbol{\beta}$. Then
$\phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta})=P(S_{\lambda}\,|\,\textbf{F}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})\times P(I_{\lambda}\,|\,\textbf{F},\boldsymbol{z})=\phi_{\lambda}(\textbf{X}\,|\,-\boldsymbol{\beta})\ ,$
and an $\textbf{X}^{*}$ optimal for $\phi_{\lambda}$ under $\boldsymbol{\beta}$ is also optimal for $\phi_{\lambda}$ under $-\boldsymbol{\beta}$.
Lemma 1 establishes events $S_{\lambda}$ and $I_{\lambda}$ are independent and provides a design equivalence result. The following equivalence theorem is more general and further simplifies the implementation of our framework:
Theorem 1.
Let $\textbf{X}^{*}$ be a locally optimal design for $\phi_{\lambda}$ under a $\boldsymbol{\beta}$ where $\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1}$. Consider an alternative $\tilde{\boldsymbol{\beta}}=\tilde{\textbf{Z}}\boldsymbol{\beta}$ where $\tilde{\textbf{Z}}$ is any diagonal matrix comprised of ±$\pm 1$. Then $\phi_{\lambda}(\textbf{X}\,|\,\tilde{\boldsymbol{\beta}})=\phi_{\lambda}(\textbf{X}\tilde{\textbf{Z}}\,|\,\boldsymbol{\beta})$ and $\tilde{\textbf{X}^{*}}=\textbf{X}^{*}\tilde{\textbf{Z}}$ is locally optimal for $\phi_{\lambda}$ under $\tilde{\boldsymbol{\beta}}.$
A consequence of Theorem 1 is that, when discussing criteria involving a known sign vector, we can assume $\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1}$ without loss of generality.
The following criterion treats $\boldsymbol{z}_{\mathcal{A}}$ as an unknown quantity:
$\phi_{\lambda}^{\pm}(\textbf{X}\,|\,\boldsymbol{\beta})=\frac{1}{2^{k}}\sum_{\tilde{z}\in\mathcal{Z}_{\mathcal{A}}}\phi_{\lambda}(\textbf{X}\,|\,\tilde{\textbf{Z}}\boldsymbol{\beta})=\frac{1}{2^{k}}\sum_{\tilde{z}\in\mathcal{Z}_{\mathcal{A}}}\phi_{\lambda}(\textbf{X}\tilde{\textbf{Z}}\,|\,\boldsymbol{\beta})\ .$ (10)
It is the expected probability of sign recovery assuming all $\tilde{\boldsymbol{z}}\in\mathcal{Z}_{\mathcal{A}}$ are equally likely. Calculating all $2^{k}$ probabilities can be computationally intensive, but Lemma 1 allows us to halve the number of computations. We state this as a corollary:
Corollary 1.
For a fixed $\mathcal{A}$ where $|\mathcal{A}|=k$, let $\mathcal{Z}_{\mathcal{A}}^{\pm}$ denote the subset of $\mathcal{Z}_{\mathcal{A}}$ including all $2^{k-1}$ unique $\boldsymbol{z}$ up to reflection. Then $\phi_{\lambda}^{\pm}(\textbf{X}\,|\,\boldsymbol{\beta})=2^{-(k-1)}\sum_{\tilde{z}\in\mathcal{Z}_{\mathcal{A}}^{\pm}}\phi_{\lambda}(\textbf{X}\tilde{\textbf{Z}}\,|\,\boldsymbol{\beta})$.
We now investigate designs that maximize $\phi_{\lambda}$ and $\phi_{\lambda}^{\pm}$. Knowledge of $\mathcal{A}$ when constructing an optimal design under either criteria leads to a trivial construction for the columns $\textbf{X}_{\mathcal{I}}^{*}$.
Proposition 1.
For a known $\mathcal{A}$, there exists a local optimal design $\textbf{X}^{*}$ where $(\textbf{I}-\textbf{P}_{1})\textbf{X}^{*}_{\mathcal{I}}=\boldsymbol{0}$, making $\phi_{\lambda}(\textbf{X}^{*}\,|\,\boldsymbol{\beta})=P(S_{\lambda}\,|\,\textbf{F}_{\mathcal{A}},\boldsymbol{\beta}_{\mathcal{A}})$ and $\phi_{\lambda}^{\pm}(\textbf{X}^{*}\,|\,\boldsymbol{\beta})=2^{-(k-1)}\sum_{\tilde{z}\in\mathcal{Z}^{\pm}_{\mathcal{A}}}P(S_{\lambda}\,|\,\textbf{F}_{\mathcal{A}}\tilde{\textbf{Z}},\boldsymbol{\beta}_{\mathcal{A}})$.
Local optimal designs exploit knowledge about $\mathcal{A}$ by completely confounding the columns of the inactive factors with the intercept, making $P(I_{\lambda}\,|\,\textbf{F},\,\boldsymbol{\beta})=1$. Finding the local optimal design, which was originally supersaturated, then reduces to finding a full column rank matrix $\textbf{X}_{\mathcal{A}}$ that maximizes probability of the $S_{\lambda}$ event(s).
To demonstrate this new theory, we will now show that if $\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1}$ is assumed known, an orthogonal design may not maximize $P(S_{\lambda}\,|\,\textbf{F}_{\mathcal{A}},\boldsymbol{\beta}_{\mathcal{A}})$. For any $\textbf{X}_{\mathcal{A}}$ the bounds of integration for $S_{\lambda}$ and the mean of $\boldsymbol{u}$ are proportional to $\textbf{V}_{\mathcal{A}}^{1/2}\boldsymbol{\beta}_{\mathcal{A}}$ and $\lambda\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}$, respectively. Hence, a design that maximizes $\textbf{V}_{\mathcal{A}}^{1/2}\boldsymbol{\beta}_{\mathcal{A}}-\lambda\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}$ may maximize $P(S_{\lambda}\,|\,\textbf{F}_{\mathcal{A}},\boldsymbol{\beta}_{\mathcal{A}})$, although the covariance matrix of $\boldsymbol{u}$ also plays a role. For any orthogonal design, this difference is $\boldsymbol{\beta}_{\mathcal{A}}-\lambda\boldsymbol{1}$. We will try to find a $\textbf{X}_{\mathcal{A}}$ with larger sign recovery probability by identifying a design whose $\textbf{V}_{\mathcal{A}}^{1/2}$ and $\textbf{C}_{\mathcal{A}}$ satisfy
$(\textbf{I}-\textbf{V}_{\mathcal{A}}^{1/2})\boldsymbol{\beta}_{\mathcal{A}}\leq(\textbf{I}-\textbf{C}_{\mathcal{A}}^{-1})\lambda\boldsymbol{1}\ .$ (11)
As $(\textbf{I}-\textbf{V}_{\mathcal{A}}^{1/2})\boldsymbol{\beta}_{\mathcal{A}}\geq\boldsymbol{0}$, we need $(\textbf{I}-\textbf{C}_{\mathcal{A}}^{-1})\lambda\boldsymbol{1}\geq\boldsymbol{0}$. For even $n\geq 6$, construct $\textbf{X}_{\mathcal{A}}$ by choosing $k\leq n-1$ columns from the $n\times n$ matrix
$\left(\begin{array}[]{c|c}2\textbf{I}-\textbf{J}&-\textbf{J}\\ \hline\cr\textbf{J}&\textbf{J}-2\textbf{I}\end{array}\right)\ ,$ (14)
where all I and $\textbf{J}=\boldsymbol{1}\boldsymbol{1}^{T}$ have size $n/2\times n/2$. Then $\textbf{V}_{\mathcal{A}}=(1-4/n^{2})\textbf{I}$. Let $k_{1}$ denote the number of columns chosen from the first $n/2$ columns and $k_{2}=k-k_{1}$ be the number from the remaining columns. If either $k_{1}$ or $k_{2}$ equal 0, $\textbf{C}_{\mathcal{A}}$ will be completely symmetric with positive off-diagonal elements $c=1-4n/(n^{2}-4)$. Thus $(\textbf{I}-\textbf{C}_{\mathcal{A}}^{-1})\lambda\boldsymbol{1}\geq\boldsymbol{0}$ and (11) becomes
$\left(1-\frac{\sqrt{n^{2}-4}}{n}\right)\boldsymbol{\beta}_{\mathcal{A}}\leq\lambda\left(1-\frac{n^{2}-4}{kn^{2}-4k(n+1)+4n}\right)\boldsymbol{1}\ .$
When $k_{1}\geq 1$ and $k_{2}\geq 1$, $\textbf{C}_{\mathcal{A}}$ and $\textbf{C}_{\mathcal{A}}^{-1}$ have a $2\times 2$ block partitioned form with completely symmetric, diagonal block matrices and constant off-diagonal block matrices. Then $\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}=(\xi_{1}\boldsymbol{1}_{k_{1}}^{T},\xi_{2}\boldsymbol{1}_{k_{2}}^{T})^{T}=\boldsymbol{\xi}$, and $\textbf{C}_{\mathcal{A}}\boldsymbol{\xi}=\boldsymbol{1}$, which gives the equations
11 11 $=\rho_{11}\xi_{1}+\rho_{12}\xi_{2}$
11 11 $=\rho_{21}\xi_{1}+\rho_{22}\xi_{2}\ ,$
where $\rho_{ij}>0$ is the unique row sum for the corresponding $k_{i}\times k_{j}$ block matrix of $\textbf{C}_{\mathcal{A}}$. Defining $\tilde{k}_{i}=(n-2k_{i})\geq 0$, we have
$\xi_{1}=\frac{\tilde{k}_{2}(n^{2}-4)}{(n-2)^{2}(k_{1}\tilde{k}_{2}+\tilde{k}_{1}k_{2})+\tilde{k}_{1}\tilde{k}_{2}}\geq 0\ ,$ (15)
with equality if and only if $k_{2}=n/2$. A similar expression holds for $\xi_{2}$ with $\tilde{k}_{1}$ in the numerator of (15), and $\xi_{2}=0$ if and only if $k_{1}=n/2$. Since the diagonal elements of $\textbf{C}_{\mathcal{A}}$ are 1, $\rho_{ii}\geq 1$ and it follows $\boldsymbol{1}-\boldsymbol{\xi}>\boldsymbol{0}$. Hence designs constructed in this way satisfy (11) for some combinations of $\boldsymbol{\beta}_{\mathcal{A}}$ and $\lambda$, but whether this implies higher sign recovery probability depends on the covariance matrix of $\boldsymbol{u}$.
For example, suppose $n=16$ and we have $p=20$ factors but we know the first $k=8$ factors are nonzero. We considered three SSDs that set all the inactive factors columns to $+1$. The first SSD set the $8$ active factor columns to an orthogonal $\textbf{X}_{\mathcal{A}}$ and the other two SSDs set active effects columns according to (14) with $k_{1}=4$ and $8$, respectively. For all designs, $\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}=\xi\boldsymbol{1}$ where $\xi=1$, $0.1575$, and $0.1607$ for the orthogonal, $k_{1}=4$, and $k_{1}=8$ designs, respectively. For the $k_{1}=4$ and $k_{1}=8$ designs, (11) is equivalent to $\lambda^{-1}\boldsymbol{\beta}_{\mathcal{A}}\leq 53.92\times\boldsymbol{1}$ and $\lambda^{-1}\boldsymbol{\beta}_{\mathcal{A}}\leq 53.72\times\boldsymbol{1}$, respectively. The three designs’ $\phi_{\lambda}$ and $\phi_{\lambda}^{\pm}$ values were compared across a range of $\lambda$ and three scenarios for $\boldsymbol{\beta}_{\mathcal{A}}$: $(0.3,0.4,\dots,1)^{T}$, $\boldsymbol{1}$, and $3\times\boldsymbol{1}$. The results for $\phi_{\lambda}$ are shown in the left panels of Figure 1.
Figure 1: Probability of sign recovery for $n=16$, $p=20$, and $k=8$ for an orthogonal design and two designs constructed from selecting columns from (14) with $k_{1}=4$ and $8$. The left and right panels correspond to $\phi_{\lambda}$ and $\phi_{\lambda}^{\pm}$, respectively.
For all scenarios, there is a range of large $\lambda$ values where all designs have $\phi_{\lambda}=0$, a middle range where the two proposed designs outperform the orthogonal design, and a range of small $\lambda$ values where the orthogonal design is superior. The orthogonal design improves over the other two designs well before condition (11) is violated, due to the role of the covariance matrices. The improvement is negligible when $\boldsymbol{\beta}_{\mathcal{A}}=3\times\boldsymbol{1}$, implying the covariance matrix is less important as the elements of $\boldsymbol{\beta}_{\mathcal{A}}$ increases. The results for $\phi_{\lambda}^{\pm}$ are shown in the right panels of Figure 1, and clearly favor the orthogonal design.
2.2 Relaxed local criteria
To make the criteria more practical, we introduce relaxations regarding the choice of $\lambda$ and the assumptions of the underlying model. To reduce the dependency on $\lambda$, we propose two summary measures of $\phi_{\lambda}$ and $\phi_{\lambda}^{\pm}$. The first summary takes the maximum sign recovery probability: $\phi_{\text{max}}(\textbf{X}\,|\,\boldsymbol{\beta})=\max_{\lambda>0}\phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta})$. However, in practice one will perform some tuning parameter selection strategy to choose a $\lambda$, and so the benefits of a design that optimizes $\phi_{\text{max}}$ will only be realized if the strategy reliably picks the corresponding $\lambda$ that maximizes $\phi_{\lambda}$. Hence, we would like a design that maintains large probabilities for a wide range of $\lambda$. It is common to calculate the lasso solution path with respect to $\log(\lambda)$. Such a transformation stretches the region of $\lambda\in(0,1)$, whose corresponding lasso estimates would receive the least amount of penalization. Therefore, we propose the criterion
[…]
The definitions for $\phi_{\max}^{\pm}(\textbf{X}\,|\,\boldsymbol{\beta})$ and $\phi_{\Lambda}^{\pm}(\textbf{X}\,|\,\boldsymbol{\beta})$ relax assumptions about $\boldsymbol{z}_{\mathcal{A}}$ and measure the maximum sign recovery and the integrated average sign recovery.
Turning to relaxing assumptions about the $\boldsymbol{\beta}$, we first assume $|\boldsymbol{\beta}_{\mathcal{A}}|\geq\beta\boldsymbol{1}$ where $\beta$ is the minimum practically significant effect size. Such a $\beta$ value is commonly chosen while performing power calculations so we will calculate our support recovery probability under the most challenging scenario of $|\boldsymbol{\beta}_{\mathcal{A}}|=\beta\boldsymbol{1}$. We will hold a discussion about incorporating uncertainty about $|\boldsymbol{\beta}_{\mathcal{A}}|$ to Section 6. To address uncertainty about $\mathcal{A}$, we fix $k\leq n-1$ and assume all supports of this size are equally likely. Denote this set of supports by $\mathcal{A}_{k}$.
With the above relaxed assumptions, we first introduce a criterion that incorporates uncertainty about $\mathcal{A}$, but treats $\boldsymbol{z}_{\mathcal{A}}$ as known. This resembles $\boldsymbol{\beta}$’s where the $Var(s+)$-optimal designs performed best in previous simulation studies. Suppose $z_{j}$ depends only on whether $j\in\mathcal{A}$ or $\mathcal{I}$, and let $\boldsymbol{z}^{*}$ be the $p\times 1$ vector comprised of the elements $z_{j}$ assuming $j\in\mathcal{A}$. Following Theorem 1, we assume without loss of generality that $\boldsymbol{z}^{*}=\boldsymbol{1}$ to identify the optimal design and then multiply the columns of the design by their actual corresponding $z_{j}^{*}$. For a given $\mathcal{A}$, let A be the $p\times p$ diagonal matrix of all zeroes except for the diagonal elements corresponding to $\mathcal{A}$, which are set to 1. Then the true sign vector is $\boldsymbol{z}=\textbf{A}\boldsymbol{1}$ and the sign-dependent criterion for a given $\beta$ is
$\Phi_{\lambda}(\textbf{X}\,|\,k,\beta)=\binom{p}{k}^{-1}\sum_{\mathcal{A}\in\mathcal{A}_{k}}\phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta}=\beta\textbf{A}\boldsymbol{1})\ ,$
being the average probability across $\mathcal{A}_{k}$ for a fixed $\lambda$. The notation $\Phi$ is intended to reflect consideration of all $\mathcal{A}\in\mathcal{A}_{k}$. The sign-independent criterion, $\Phi_{\lambda}^{\pm}(\textbf{X}\,|\,k,\beta)$, is similarly defined, considering all $2^{k}$ sign vectors as equally likely. The two summary measures originally defined on $\phi_{\lambda}$ and $\phi_{\lambda}^{\pm}$ are straightforward to define on these new criteria and share similar notation, i.e., $\Phi_{\text{max}}(\textbf{X}\,|\,k,\beta)$ and $\Phi_{\Lambda}(\textbf{X}\,|\,k,\beta)$. Computational details of the criteria may be found in the Supplementary Materials.
Optimizing any of the $\Phi$-type criteria is both analytically and algorithmically challenging. Indeed, simply evaluating such criteria for a given X can be cumbersome. We next discuss properties of these criteria and, motivated by approximate design theory (Pukelsheim, 2006), we propose a more tractable framework centered on finding optimal C matrices. Section 4 leverages the optimal forms of C to identify optimal, or at least nearly-optimal, exact designs.
3 Optimal Correlation Matrices
The criteria from Section 2.2 assume $|\boldsymbol{\beta}_{\mathcal{A}}|=\beta\boldsymbol{1}$ and evaluate an X across all possible supports, and so are invariant to column permutations of X. Additionally, $\Phi^{\pm}_{\lambda}$ and its summaries across $\lambda$ are invariant to sign-flipping the columns of X. Traditional eigenvalue-based criteria, being functions of $\textbf{M}=\textbf{X}^{T}(\textbf{I}-\textbf{P}_{1})\textbf{X}$, also enjoy these two invariance properties, expressed as simultaneous row/column permutations and sign transformations of M. Indeed, these properties have been exploited to find optimal forms of M when the domain of the functions is expanded to all symmetric, positive definite matrices, denoted $\mathcal{M}$. For an $\textbf{M}\in\mathcal{M}$, define its permutation-averaged form to be $\overline{\textbf{M}}=(p!)^{-1}\sum_{\boldsymbol{\Pi}\in\mathcal{P}}\boldsymbol{\Pi}\,\textbf{M}\,\boldsymbol{\Pi}^{T}$ where $\mathcal{P}$ is the set of all $p\times p$ permutation matrices. Then $\overline{\textbf{M}}\in\mathcal{M}$ is a compound symmetric matrix. Further averaging of $\overline{\textbf{M}}$ across all $2^{p}$ sign transformations leads to a matrix proportional to the identity matrix, i.e., an M for an orthogonal design. Assuming concavity, the criterion value for $\overline{\textbf{M}}$ is greater than or equal to that for M. Hence the search for the optimum M may be restricted to the class of compound symmetric matrices in $\mathcal{M}$.
[…]
The probabilities for events (7) and (8) for $\textbf{C}\in\mathcal{C}$ are given in the following lemma:
Lemma 2.
[…]
$\boldsymbol{u}$ $\sim N\left(\frac{\lambda\sqrt{n}}{1-c}\left[\boldsymbol{1}-z_{\mathcal{A}}\gamma\boldsymbol{z}_{\mathcal{A}}\right],\ \frac{1}{1-c}\left[\textbf{I}_{k}-\gamma\boldsymbol{z}_{\mathcal{A}}\boldsymbol{z}_{\mathcal{A}}^{T}\right]\right)$
$\boldsymbol{v}$ $\sim N\left(\lambda\sqrt{n}z_{\mathcal{A}}\gamma\boldsymbol{1},(1-c)\left[\textbf{I}+\gamma\textbf{J}\right]\right)$
with $z_{\mathcal{A}}=\boldsymbol{1}^{T}\boldsymbol{z}_{\mathcal{A}}$ and $\gamma=c/(1+c(k-1))$. Moreover, if $\boldsymbol{\beta}_{\mathcal{A}}$ does not depend on $\mathcal{A}$, the probabilities of the two events are constant across all such $\mathcal{A}$.
The criterion $\Phi_{\lambda}(\textbf{X}\,|\,k,\beta)$ meets the latter condition of Lemma 2 and assumes $\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1}$, leading to the random vectors
$\boldsymbol{u}\sim N\left(\frac{\lambda\sqrt{n}}{1+c(k-1)}\boldsymbol{1},\frac{1}{1-c}\left[\textbf{I}-\gamma\textbf{J}\right]\right),\qquad\boldsymbol{v}\sim N\left(\lambda\sqrt{n}k\gamma\boldsymbol{1},(1-c)\left[\textbf{I}+\gamma\textbf{J}\right]\right)\ .$
The corresponding criterion defined with respect to $\mathcal{C}$ and for a fixed $\lambda$ is denoted
$\psi_{\lambda}(c\,|\,k,\beta)=P(S_{\lambda}\,|\,c,\boldsymbol{\beta}_{\mathcal{A}}=\beta\boldsymbol{1})\times P(I_{\lambda}\,|\,c,\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1})\ .$ (16)
The analog to $\Phi_{\lambda}^{\pm}(\textbf{X}\,|\,k,\beta)$ is denoted
$\psi_{\lambda}^{\pm}(c\,|\,k,\beta)=\frac{1}{2^{k-1}}\sum_{\tilde{\boldsymbol{z}}\in\mathcal{Z}_{\mathcal{A}}^{\pm}}P(S_{\lambda}\,|\,c,\boldsymbol{\beta}_{\mathcal{A}}=\beta\tilde{\boldsymbol{z}})\times P(I_{\lambda}\,|\,c,\boldsymbol{z}_{\mathcal{A}}=\tilde{\boldsymbol{z}})\ .$ (17)
The summarized versions of (16) and (17) across $\lambda$ will be denoted similarly to their exact design counterparts from Section 2, replacing $\Phi$ with $\psi$.
The criteria $\psi_{\lambda}$ and $\psi_{\lambda}^{\pm}$ involve a single variable, but are still challenging to optimize analytically. Numerical optimization of these criteria, however, is straightforward and computationally efficient. We demonstrate numerical optimization of these new criteria for the situation of $p=10$, $k=4$, $\beta=2$, and $n=10$. Figure 2 shows the contour plots of $\psi_{\lambda}$ and $\psi_{\lambda}^{\pm}$. The optimal $c$ values for the corresponding $\psi_{\Lambda}$ and $\psi_{\Lambda}^{\pm}$ are $0.14$ and 00, respectively. The resulting optimal C matrices then match the ideal forms for the $Var(s+)$ and $UE(s^{2})$-criterion. While these ideal C are not possible for SSDs, this example provides further justification for these two heuristic SSD criteria for sign recovery under the lasso.
Figure 2: Contour plots of sign recovery probabilities when C is completely symmetric across possible values of the off-diagonal $c$ and $\log(\lambda)$ where $n=10$ and $k=4$ factors are active with $\beta=2$.
[…]
We now study the behavior of $\psi_{\lambda}$ and $\psi^{\pm}_{\lambda}$ in a neighborhood about $c=0$. Direct application of the multivariate Leibniz integral rule gives the following lemma.
Lemma 3.
For all $\mathcal{A}$ where $|\mathcal{A}|=k$ and $\beta>0$, $\frac{d}{dc}P(I_{\lambda}|c,\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1})\big{|}_{c=0}=0$ for all $\lambda$ and $\frac{d}{dc}P(S_{\lambda}|c,\boldsymbol{\beta}_{\mathcal{A}}=\beta\boldsymbol{1})\big{|}_{c=0}\geq 0$ for $\lambda$ satisfying
$2\lambda_{n}$ $\geq\frac{g(\tau_{n})}{G(\tau_{n})}\ ,$ (18)
where $G(\cdot)$ and $g(\cdot)$ represent the $N(0,1)$ cumulative distribution function and probability density function, respectively, with $\lambda_{n}=\lambda\sqrt{n}$ and $\tau_{n}=\sqrt{n}(\beta-\lambda)$.
A direct consequence of Lemma 3 is that $\psi_{\lambda}(0|k,\beta)$ can be improved upon by some $c>0$. We state this important result as a theorem.
Theorem 2.
For all $\mathcal{A}$ where $|\mathcal{A}|=k$, $\frac{d}{dc}\psi_{\lambda}(c|k,\beta)\big{|}_{c=0}>0$ for $\lambda$ satisfying (18). Hence there exists some $c_{\lambda}>0$, where $\psi_{\lambda}(c_{\lambda}|k,\beta)>\psi_{\lambda}(0|k,\beta)$.
Applying this result to the situation shown in Figure 2 with known sign, inequality (18) holds for $\log(\lambda)\geq-22.763$, which covers the entire region of $\lambda$ values for which $\psi_{\lambda}(c)>0$. It follows then that some $c>0$ will maximize the summary measures $\psi_{\Lambda}(c)$ or $\psi_{\max}(c)$. Theorem 2 thus provides a mathematically rigorous justification that the ideal designs under the $Var(s+)$-criterion maximize the probability of sign recovery for known sign.
The following theorem establishes a similar justification for heuristic orthogonality measures under unknown signs.
Theorem 3.
For all $\mathcal{A}$ where $|\mathcal{A}|=k$ and $\beta>0$, $\frac{d}{dc}\psi_{\lambda}^{\pm}(c|k,\beta)\big{|}_{c=0}=0$ for all $\lambda$. Moreover, $c=0$ is a local maximum for $\psi_{\lambda}^{\pm}(c|k,\beta)$ when
$\begin{split}\frac{q}{\binom{k}{2}}\ \frac{\lambda_{n}g(\lambda_{n})}{G(\Delta_{n})}&\left(k(1-\lambda_{n}^{2})+(q-1)\frac{\lambda_{n}g(\lambda_{n})}{G(\Delta_{n})}\right)\leq\\ &\frac{g(\tau_{n})}{G(\tau_{n})}\left(\beta_{n}+\lambda_{n}+\lambda_{n}^{2}\tau_{n}-\left[\frac{\beta_{n}^{2}-\lambda_{n}^{2}}{2}+\beta_{n}\lambda_{n}\right]\frac{g(\tau_{n})}{G(\tau_{n})}\right)\end{split}$ (19)
where $\beta_{n}=\beta\sqrt{n}$ and $G(\Delta_{n})=G(\lambda_{n})-G(-\lambda_{n})$.
Applying this result to the situation shown in Figure 2 with unknown sign, inequality (69) holds for $\log(\lambda)\in[-0.988,0.640]$. For $\log(\lambda)$ outside this region, there are clearly some $c>0$ for which $\psi_{\lambda}^{\pm}(c)>0$, although the probabilities are relatively small. These small probabilities do not influence $\psi_{\max}^{\pm}$ and have minimal influence on $\psi_{\Lambda}^{\pm}$, making $c=0$ a global maximum for both criteria. We conjecture that generally $c=0$ and some $c>0$ are global maxima for the unknown and known sign criteria, respectively.
4 Fast Design Construction with HILS
Section 3 identified optimal forms of completely symmetric C matrices under different assumptions about $\boldsymbol{z}$. Unfortunately, no SSD exists whose C achieves these forms because $n<p+1$. Ideally, one would implement a design search algorithm that ranks SSDs according to a summary criterion of the $\Phi_{\lambda}$- or $\Phi_{\lambda}^{\pm}$-criterion. However, these criteria demand intense computations to evaluate a single SSD and search algorithms require many evaluations. We now describe an algorithmic construction that compromises the rigorous but computationally-intensive criteria in Section 2 with new computationally-efficient heuristic criteria based on the theory in Section 3.
[…]
To construct designs based on multiple criteria, Cao et al. (2017) combined the coordinate exchange algorithm (Meyer and Nachtsheim, 1995) with an elitist-like operator to produce $ncand$ designs lying near the Pareto front of the criteria. A set of $ntry$ initial designs are provided and each will generate its own Pareto set. That is, for each initial design, many exploratory designs are generated via coordinate exchanges. Only those exploratory designs that either dominate the initial design, or are non-dominated by the other exploratory designs form the initial design’s Pareto set. The procedure continues until each non-dominated design cannot be further improved via coordinate exchange. The approximate Pareto fronts from each initial design are then combined into a single front comprised of the non-dominated designs across all $ntry$ Pareto sets, giving a set of, say, $ncand$ designs. Once the final approximate Pareto set is constructed, we identify which of its $ncand$ designs maximizes the lasso criterion of interest.
[…]
Our proposed construction algorithm first identifies a set of $ncand$ candidate designs through the Pareto front optimization search algorithm from Cao et al. (2017) based on our two lasso-motivated criteria. We then employ our exact lasso sign recovery criteria from Section 2 as a secondary criterion, as it more accurately measures a design’s screening capabilities, to rank the $ncand$ designs and choose the best one. We refer to this algorithm as the Heuristic-Initiated Lasso Sieve, or HILS.
[…]
Figure 3: Plot of the approximate Pareto front for the three designs selected by the coordinate exchange algorithm for the $n=9$, $p=10$, unknown signs case. The starting criteria values for each of the $ntry=20$ starting designs from the algorithm are plotted in black. Each point represents a different design. The criteria values for the DCD are also plotted to show that the approximate Pareto Front designs the DCD in terms of the two criteria.
Although HILS can significantly reduce the number of $\Phi_{\lambda}$ evaluations compared to constructing an optimal design using, for instance, a coordinate exchange algorithm, there are settings when even a more modest number of evaluations is not computationally feasible. The same issue was encountered by Singh and Stufken (2022). We now discuss some ways to efficiently evaluate the $\Phi_{\lambda}$ and $\Phi_{\lambda}^{\pm}$-criteria, and hence their corresponding summarized versions. For a support size of $k$ and design X, one full evaluation of $\Phi_{\lambda}$ and $\Phi_{\lambda}^{\pm}$ requires consideration of all $\binom{p}{k}$ supports of size $k$. For $\Phi_{\lambda}^{\pm}$, an additional $2^{k}$ computations are required for each $\mathcal{A}$ to consider all sign vectors. Evaluating either criteria would benefit from a reduction in the number of supports considered. Sampling from $\mathcal{A}_{k}$ completely at random is intuitive, but may require a large number of samples to be representative. As high correlations between two factors can result in diminished lasso support recovery performance due to low $P(I_{\lambda}\,|\,\textbf{F},\,\boldsymbol{z})$, a support sampling method that balances (or approximately balances) the number of times pairwise sets of factors are included in $\mathcal{A}$ together can be advantageous. Smucker and Drew (2015) utilized nearly balanced incomplete block designs (NBIBDs) to approximate a model space with only a relatively small number of blocks, in the context of model-robust optimal designs. Hence, we recommend implementing the NBIBD sampling method to adequately represent $\mathcal{A}_{k}$ with between 64-128 supports for modest $p$ and $k$ values. We denote such a subset of supports by $\tilde{\mathcal{A}}_{k}$.
From Lemma 1, for a fixed $\mathcal{A}$, the probabilities are equal between $\boldsymbol{z}_{\mathcal{A}}$ and $-\boldsymbol{z}_{\mathcal{A}}$. Thus, only $2^{k-1}$ sign vectors need be considered for the $\Phi_{\lambda}^{\pm}$-criterion. While this cuts the computation in half, prior knowledge from the practitioner can be leveraged to select an even smaller set of representative sign vectors to evaluate. For example, if there is no knowledge on sign direction, the most likely sign vectors are those with an equal, or nearly equal, number of ±$\pm 1$. For $k=10$, one would need to only consider, up to reflection, the $126$ possible sign vectors having five $+1$’s and $-1$’s rather than all $5122^{10-1}=$ 512 vectors. There may also be strong prior belief about the signs of some or all of the factors that can reduce the set of sign vectors to a manageable number, or even to a single element.
5 Examples
We now demonstrate near-optimal SSDs found under our criteria via HILS under three example scenarios. For each case, the coordinate exchange algorithm was executed with $ntry=50$ random starting designs and identified $ncand$ final designs on the approximate Pareto front of the criteria described in Section 4. Of these $ncand$ designs, the final HILS design was selected using the $\Phi_{\text{max}}^{\pm}$ criterion. For each scenario, we provide the computational time that it took to find the HILS-optimal design on a Desktop with an Intel ® Core i7-6700 CPU @ 3.40GHz processor with 16GB RAM. We then compare the best identified designs by inspecting the $\Phi_{\lambda}$ and $\Phi_{\lambda}^{\pm}$ curves as a function of $\log(\lambda)$. These curves represent the average probability of sign recovery of the lasso solution path over a range of tuning parameter values. In comparing designs in this way, we eliminate the need to choose a tuning parameter. The design with larger $\Phi_{\lambda}$ or $\Phi_{\lambda}^{\pm}$ over a larger range of $\log(\lambda)$ is considered to be the best performing.
Additionally, we evaluate the performance of the HILS designs compared to competitors under $k$ and $\beta$ settings that are different to the $k$ and $\beta$ values used to construct the HILS designs. This evaluates the sensitivity of the HILS construction method to $k$ and $\beta$.
5.1 Scenario 1: $n=9$, $p=10$ Unknown Signs
The HILS algorithm was performed for $k=3$ and $\beta=3$ assuming unknown signs. The HILS-optimal design was selected in approximately six minutes of computation time. Because $n\approx p$, this scenario is close to one in which an orthogonal design exists. Therefore, this represents an experiment in which other optimal design criteria should perform rather well. For unknown signs, the optimal $c$ value is taken as $c^{*}=0$ following Theorem 3. Since $\textbf{C}^{*}$ is orthogonal and the $UE(s^{2})$ criterion targets orthogonality directly, the $UE(s^{2})$-optimal designs were deemed appropriate for comparison. Additionally, since the two criteria used to construct the DCDs are optimized under an orthogonal design, the DCD was also compared.
Figure 4 compares the $\Phi_{\lambda}^{\pm}$ curves as a function of $\log(\lambda)$ for the HILS-optimal design, DCD (Singh and Stufken, 2022), and a $UE(s^{2})$-optimal design under a variety of $k$ and $\beta$ settings. The top-middle plot represents the $k$ and $\beta$ setting in which the HILS design was optimized for. The top row of Figure 4, representing $k=3$, demonstrates that the HILS design out performs the others in terms of lasso sign recovery regardless of the $\beta$ value. When $k=5$, Figure 4 shows little difference between the 3 designs. The HILS and $UE(s^{2})$- optimal designs show a larger $\Phi_{\lambda}^{\pm}$ for larger $\log(\lambda)$, but the DCD attains a slightly larger maximum of $\Phi_{\lambda}^{\pm}$ (for $\beta=3$ and $\beta=4$). Overall, the HILS-optimal design saw better performance than its competitors under the $k$ value that it was optimized for and comparable performance as the $k$ assumption was violated.
Figure 4: $\Phi_{\lambda}^{\pm}$ curves as a function of $\log(\lambda)$ for the HILS-optimal design, DCD, and a $UE(s^{2})$-optimal design under a variety of $k$ and $\beta$ settings for scenario 1.
5.2 Scenario 2: $n=9$, $p=10$ Known Signs
To build on the $n$ and $p$ settings in scenario 1, the HILS algorithm was then performed for $k=3$ and $\beta=3$ assuming known signs. Here, the optimal $c$ value was found via numerical optimization to be taken as $c^{*}\approx 0.04$. The HILS algorithm selected the HILS-optimal design in approximately 4 minutes.
Figure 5 compares the $\Phi_{\lambda}$ curves as a function of $\log(\lambda)$ for the HILS-optimal design, DCD (Singh and Stufken, 2022), and a $Var(s+)$-optimal design under a variety of $k$ and $\beta$ settings. To demonstrate the added value these three designs have by incorporating sign information, the figure also has the $\Phi_{\lambda}$ curve for a $UE(s^{2})$-optimal design. The top-middle plot represents the $k$ and $\beta$ setting in which the HILS design was optimized for. For all $\beta$ and $k$ settings except $\beta=4$ and $k=5$, the HILS-optimal design shows a larger $\Phi_{\text{max}}$ and area under the curve compared to DCD and $Var(s+)$. For $\beta=4$ and $k=5$, Figure 5 shows a slightly larger $\Phi_{\text{max}}$ for DCD compared to the HILS-optimal design. However, the HILS-optimal design sees higher $\Phi_{\lambda}$ for larger values of $\log(\lambda)$ (around $\log(\lambda)=1$) and has a larger area under its $\Phi_{\lambda}$ curve. Thus, we deem that the HILS-optimal design is superior to the competitors in all settings demonstrated in Figure 5. Finally, we note the poor performance of the $UE(s^{2})$-optimal design compared to all designs, especially when $k=5$ with $\Phi_{\lambda}\approx 0$.
Figure 5: $\Phi_{\lambda}$ curves as a function of $\log(\lambda)$ for the HILS-optimal design, DCD, and a $Var(s+)$-optimal design under a variety of $k$ and $\beta$ settings under scenario 2.
5.3 Scenario 3: $n=14$, $p=20$ Known Signs
This scenario assumed known effect directions with $k=5$ and $\beta=2$. The optimal $c$ value was found via numerical optimization to be taken as $c^{*}\approx 0.045$. With a larger difference between $n$ and $p$, the Pareto front found by the coordinate exchange algorithm consisted of 14 designs (the Pareto fronts consisted of 3 designs for Scenarios 1 and 2). The HILS algorithm selected the HILS-optimal design in approximately 1 hour and 21 minutes.
Figure 6 compares the $\Phi_{\lambda}$ curves as a function of $\log(\lambda)$ for the HILS-optimal design, DCD (Singh and Stufken, 2022), and a $Var(s+)$-optimal design under a variety of $k$ and $\beta$ settings. We again included the $\Phi_{\lambda}$ curve for a $UE(s^{2})$-optimal design. The top-middle plot represents the $k$ and $\beta$ setting in which the HILS design was optimized for. Across all $\beta$ and $k$ settings there is very little difference between the performances of the three designs that include sign information. However, all of these designs had significantly higher $\Phi_{\lambda}$ values than the $UE(s^{2})$-optimal design.
Figure 6: $\Phi_{\lambda}$ curves as a function of $\log(\lambda)$ for the HILS-optimal design, DCD, and a $Var(s+)$-optimal design under a variety of $k$ and $\beta$ settings under scenario 3.
The lack of separation between the HILS-optimal design from the other designs is not unexpected. The larger the discrepancy between $p$ and $n$, the less we expect $||\textbf{C}-\textbf{C}^{*}||_{F}$ to be a good measure because we are dealing with a scenario in which C will never be close to compound symmetric. In scenarios 1 and 2, $n$ is close to $p$, and, thus, C can be close to the ideal structure of compound symmetry. Nonetheless, the HILS-optimal design was found much more efficiently than the DCD which took on the order of dozens of hours. Since the HILS-optimal design can be constructed more quickly and show comparable performance to the DCDs, we view this algorithmic design construction approach as worthwhile.
6 Summary
The SSD literature has predominately constructed designs by optimizing heuristic criteria that measure a design’s proximity to a (nonexistent) orthogonal design. The heuristic criteria are tractable in their optimization, in that optimal designs can generally be constructed directly or algorithmically in a reasonable amount of time. However, these heuristic criteria are not directly tied to a screening analysis method, so there is no guarantee that the resulting analysis under an optimal design will have good statistical properties. This article resolves this disconnect by optimizing criteria based on the probability of sign recovery under the lasso, which is well-defined even when $n-1<p$. Our major contributions are:
  1. 1.
    A local optimality criterion assuming known $\boldsymbol{\beta}$ and fixed $\lambda$. A trivial design that confounds all inactive factors with the intercept is shown to be optimal for sign recovery. An exact design construction is given that can improve the probability of sign recovery over an orthogonal design for some $\lambda$ The design has positive and nearly constant pairwise correlations, following the ideal structure for the $Var(s+)$-criterion.
  2. 2.
    More practical criteria that relax the assumptions about $\boldsymbol{\beta}$ and $\lambda$. Such criteria are computationally intensive and hence difficult to optimize both analytically and algorithmically, requiring computations across all supports of a given size and potentially many sign vectors.
  3. 3.
    A study of the optimal form of C with and without known $\boldsymbol{z}$. By conditioning on completely symmetric matrices, we arrive at a univariate optimization problem. The framework mimics the approximate design approach for least-squares analyses.
  4. 4.
    A rigorous proof that across all completely symmetric matrices, an orthogonal correlation matrix is suboptimal for nearly all $\lambda>0$ when $\boldsymbol{z}$ is known. The optimal form instead takes on positive, constant off-diagonal elements. In the case of unknown sign information, an orthogonal correlation matrix is shown to be a local maximum for a range of $\lambda$. These optimal forms rigorously justify the $UE(s^{2})$- and $Var(s+)$-criterion in the cases of unknown and known $\boldsymbol{z}$, respectively. This at least partially solves the two open problems from Section 1.1.
  5. 5.
[…]
Our work provides an alternative to using simulation to compare SSDs, which can be tedious and difficult to reproduce independently. In particular, there are at least two reasons to be skeptical of conclusions drawn from simulations in the context of SSDs. First, simulation can be misleading when subtly different versions of complicated statistical procedures are used. In our attempts to reproduce and compare the simulation studies of Singh and Stufken (2022) and Weese et al. (2021), we discovered the regularization methods used were sensitive to two different aspects of the procedure implementation. In particular, these papers used the Gauss-Dantzig selector with different ways of exploring the tuning parameter space. They also differed in how they implemented a secondary thresholding of the estimates which determines a subset of the active factors that should remain active. Typically this threshold removes estimates whose absolute magnitude are less than $\sigma^{2}$, but thresholding in this way is of dubious reliability because there is no natural way to estimate $\sigma^{2}$ and thus threshold levels are more or less arbitrary. The approach we present in this paper avoids these difficulties by eliminating the need for simulation at all, in a similar way that closed-form power analysis procedures are routinely used in simple, replicated experimental design settings.
Another danger of the existing heuristic criteria is that there is no guarantee the optimal design has any statistical value. For example, there technically exists a $UE(s^{2})$-optimal design with $n=2$ runs and $p=100$ factors, but this design has no statistical value. Our criteria, however, do reflect statistical value and so can be used to give objective information about design quality, akin to a power analysis. Experimenters can specify a minimum effect size of interest, $\beta$, as well as an educated guess regarding the number of factors that are likely to have such an effect size and investigate the sign-recovery probabilities as a function of $\lambda$. If the maximum average probability is close to 0, it is unlikely that the design can provide reliable sign recovery for the specified effect size. Such a low value could lead to a reconsideration of the runs budget, expected sparsity, and/or the size of effects that are deemed of interest to detect.
Our framework allows an even deeper investigation, if desired. For example, we can determine whether a low average probability is due to an inability to reliably exclude all inactive effects (the $I_{\lambda}$ event) or to reliably include all active effects (the $S_{\lambda}$ event). In screening experiments, we are much more concerned with identifying the true effects than with allowing inactive effects through the filter. This suggests that the $S_{\lambda}$ event is more important to the experimenter than the $I_{\lambda}$ event.
There are many avenues of future research to be explored. First, it is important to establish the concavity (or log concavity) of the criteria that are summarized over $\lambda$. We have already seen that such criteria are not concave for a fixed $\lambda$. Second, new construction techniques for designs approaching the desired structure of the $Var(s+)$-criterion would be incredibly useful. Continuing with exact design construction, while our algorithm is faster relative to Singh and Stufken (2022), it is much slower than that for the heuristic criteria. More work is needed to make the design construction algorithms faster. Finally, we are currently extending the criteria to include the case where inactive factors have small, but nonzero, effects as well as considering the case of criteria for a thresholded lasso.
SUPPLEMENTARY MATERIAL
Following the References are proofs and computational details on the evaluation of the criteria. R files that construct designs following (14); identify optimal $c$ values for the criteria in Section 3; and construct designs with the HILS algorithm. may be found at https://github.com/hkyoung361/Lasso_Optimal_SSD. Code is also included for the examples in Section 5.