### v3

# An Optimal Design Framework for Lasso Sign Recovery

Abstract
Supersaturated designs investigate more factors than there are runs, and are often constructed under a criterion measuring a design’s proximity to an unattainable orthogonal design. The most popular analysis identifies active factors by inspecting the solution path of a penalized estimator, such as the lasso. Recent criteria encouraging positive correlations between factors have been shown to produce designs with more definitive solution paths so long as the active factors have positive effects. Two open problems affecting the understanding and practicality of supersaturated designs are: (1) do optimal designs under existing criteria maximize support recovery probability across an estimator’s solution path, and (2) why do designs with positively correlated columns produce more definitive solution paths when the active factors have positive sign effects? To answer these questions, we develop criteria maximizing the lasso’s sign recovery probability. We prove that an orthogonal design is an ideal structure when the signs of the active factors are unknown, and a design constant small, positive correlations is ideal when the signs are assumed known. A computationally-efficient design search algorithm is proposed that first filters through optimal designs under new heuristic criteria to select the one that maximizes the lasso sign recovery probability.
Keywords: Constrained-positive $\text{Var}(s)$-criterion; Gauss Dantzig selector; Supersaturated designs; $\text{UE}(s^{2})$-criterion; Variable selection
1 Introduction
A screening experiment aims to learn which $k$ out of $p$ factors most influence the response variable using as few runs, $n$, as possible. Achieving this goal with a traditional least-squares analysis is challenging because often not all possible factorial effects can estimated simultaneously. Still, effective screening is feasible under the effect sparsity principle and when the experimental design is carefully selected (Box and Hunter, 1961a, b; Mee, 2009; Xu et al., 2009; Mee et al., 2017). Two-level supersaturated designs, or SSDs, push screening to its limits with $n<p+1$ so that not even the $p$ main effects can be estimated simultaneously via least-squares. The recent literature on SSD analysis instead promotes penalized regression methods to perform screening under the main effects model, but the literature on the construction of SSDs has not kept up with this development. The present work addresses the gap between analysis and construction by developing an optimal design framework that tailors design selection to statistical properties of the lasso estimator.
Throughout this paper, we will assume a main effects model $\boldsymbol{y}=\beta_{0}\boldsymbol{1}+\textbf{X}\boldsymbol{\beta}+\boldsymbol{e}$ where X is the $n\times p$ design matrix with elements $x_{ij}=\pm 1$, $\boldsymbol{\beta}=(\beta_{1},\dots,\beta_{p})^{T}$ is a sparse vector with $k<p$ nonzero elements, and $\boldsymbol{e}\sim N(\boldsymbol{0},\sigma^{2}\textbf{I})$. Without loss of generality, assume $\sigma^{2}=1$, making $\boldsymbol{\beta}$ the signed signal-to-noise ratios. Then the analysis goal is estimation of the support of $\boldsymbol{\beta}$, denoted $\mathcal{A}=\{j:|\beta_{j}|>0\}$, although for technical reasons we will be more interested in estimating the sign vector of $\boldsymbol{\beta}$, denoted $\boldsymbol{z}=\text{sign}(\boldsymbol{\beta})$. We will focus our attention on optimal SSDs, although the framework developed in this paper could be applied to more general screening designs. Note that when we indicate designs are “optimal”, we are referring to the best designs found by heuristic optimality procedures, unless otherwise indicated.
1.1 Background and Open Problems
The least-squares estimator for $\boldsymbol{\beta}$ is not unique for any SSD, complicating its analysis and design construction. If there were SSDs with unique least-squares estimators, the ideal X would satisfy $\textbf{S}=\textbf{L}^{T}\textbf{L}=n\textbf{I}_{p+1}$ where $\textbf{L}=(\boldsymbol{1}|\textbf{X})$, as its $\hat{\beta}_{j}$ would have the minimum possible variance across all SSDs. These so-called orthogonal designs only exist when $n=0\,(\text{mod}\ 4)>p$. For arbitrary $n>p$, a design could instead be selected by minimizing a variance-based criterion such as the $D$-criterion, $|(\textbf{L}^{T}\textbf{L})^{-1}|$, or $A$-criterion, $\text{tr}[(\textbf{L}^{T}\textbf{L})^{-1}]$. This optimal design framework based on least-squares is well-developed and tractable (Pukelsheim, 2006; Goos and Jones, 2011), but is not applicable for SSDs.
Most SSD construction methods focus on optimizing heuristic criteria that find a design whose S is close to $n\textbf{I}_{p+1}$. Booth and Cox (1962) proposed the criterion $\text{E}(s^{2})=2/(p(p-1))\sum_{1<i<j\leq p}s_{ij}^{2}$ assuming balanced designs (i.e., $\textbf{X}^{T}\boldsymbol{1}=\boldsymbol{0}$), which has been extensively studied (see Lin, 1993; Wu, 1993; Nguyen, 1996; Li and Wu, 1997; Tang and Wu, 1997; Ryan and Bulutoglu, 2007). Jones and Majumdar (2014) proposed the unconditional $\text{E}(s^{2})$-criterion, $\text{UE}(s^{2})=2/(p(p+1))\sum_{0\leq i<j\leq p}s_{ij}^{2}$, that includes the $s_{0j}^{2}$ elements. Other criteria have been proposed that measure a design’s quality in terms of subsets of columns (Deng et al., 1996, 1999; Jones et al., 2009), which is connected to a stepwise or all-subsets analysis (Lin, 1995; Abraham et al., 1999; Westfall et al., 1998; Liu et al., 2007). Chen and Lin (1998) and Sarkar et al. (2009) investigated support recovery properties of SSDs under such sequential analyses.
Li and Lin (2002) and Phoa et al. (2009) advocated for SSDs to be analyzed under penalized estimation that induces sparse estimates. The previously mentioned SSD criteria, however, are not directly tied to statistical properties of such an analysis. There have been some attempts in the literature to address this disconnect. Marley and Woods (2010), Draguljić et al. (2014), and Weese et al. (2015) have compared optimal SSDs under different criteria via simulation with the lasso and/or Gauss-Dantzig selector (GDS), finding little difference between the SSDs in terms of support recovery. Additionally, the Dantzig selector and its statistical properties are closely related to the lasso estimator (Meinshausen et al., 2007; Lounici, 2008; Bickel et al., 2009; James et al., 2009; Asif and Romberg, 2010), but the mathematics of the lasso are more tractable, making its optimal design framework more straightforward.
Recently, Weese et al. (2017) and Weese et al. (2021) proposed $\text{Var}(s+)$-optimal designs that minimize
$\text{Var}(s)=\text{UE}(s^{2})-\text{UE}(s)^{2}\text{ such that }\frac{\text{UE}^{*}(s^{2})}{\text{UE}(s^{2})}>c\text{ and }\text{UE}(s)>0\ ,$
where $\text{UE}^{*}(s^{2})$ is the optimal $\text{UE}(s^{2})$ value for the given $n$ and $p$, and $\text{UE}(s)=2/(p(p+1))\sum_{0\leq i<j\leq p}s_{ij}$. The ideal structure of S under $\text{Var}(s+)$ is notably non-orthogonal, having average off-diagonal elements equal to a small, positive constant, making $\text{Var}(s)=0$ while satisfying the constraints on $\text{UE}(s^{2})$ and $\text{UE}(s)$. Simulation studies showed $\text{Var}(s+)$-optimal designs (with $c=0.8$) performed similarly to other optimal SSDs under the GDS, but when the effect directions were known in advance (i.e. the nonzero signs of $\boldsymbol{\beta}$ were all positive), the $\text{Var}(s+)$-optimal design had better support recovery properties. The simulation studies mentioned above only perform relative comparisons of SSDs generated by heuristic criteria and so do not answer the following open problems:
  1. 1.
    How well do existing heuristic criteria find SSDs that maximize the probability of support recovery across all possible SSDs?
  2. 2.
    Why does the $\text{Var}(s+)$-criterion identify SSDs with better support recovery probability when the active effects have positive signs?
The theory developed in this paper answers both of these problems.
Singh and Stufken (2023) addressed problem 1 by proposing two criteria related to the restricted isometry property (Candes and Tao, 2007) and irrepresentable condition (Gai et al., 2013) of the GDS, which are sufficient for estimation and sign recovery consistency. For given $\mathcal{A}$ of size $k=|\mathcal{A}|$, let $\textbf{C}_{\mathcal{A}}$ denote the correlation matrix of the columns in $\mathcal{A}$. Let $0\leq\lambda_{\mathcal{A}1}\leq 1$ and $\lambda_{\mathcal{A}k}\geq 1$ denote the minimum and maximum eigenvalues of $\textbf{C}_{\mathcal{A}}$, respectively, and define $\tau_{\mathcal{A}}=\max(1-\lambda_{\mathcal{A}1},\lambda_{\mathcal{A}k}-1)$. Singh and Stufken (2023) propose calculating the 95th percentile of all $\tau_{\mathcal{A}}$, denoted by $\zeta_{k}$. They then calculate $\zeta_{k}$ for $k=2,\dots,\lceil n/2\rceil$ and define the estimation capacity criterion:
$\text{EC}=\text{mean}(\zeta_{2},\dots,\zeta_{\lceil n/2\rceil})\ .$
Ideally, EC would be 0, corresponding to an orthogonal design. Their second criterion assumes the sign of the active effects is always positive, denoted by $\boldsymbol{1}_{k}$. For a given $\mathcal{A}$, the following two weak sign consistency conditions are checked
$|\textbf{C}_{\mathcal{I}\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}_{k}|$ $<\boldsymbol{1}_{p-k}\ ,$ (1)
$|\textbf{C}_{\mathcal{I}\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\text{sign}(\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}_{k})|$ $<\boldsymbol{1}_{p-k}\ ,$ (2)
where $\textbf{C}_{\mathcal{I}\mathcal{A}}$ is the matrix of cross correlations of columns between the inactive set, =$\mathcal{I}=\{j:\beta_{j}=0\}$, and $\mathcal{A}$. Similar to EC, this condition is checked across all $\mathcal{A}$ where $|\mathcal{A}|=k$ and $\gamma_{k}$ denotes the proportion of such $\mathcal{A}$ that achieve both (1) and (2). Their weak sign consistency criterion is defined to be
$\text{WSC}=\text{mean}(\gamma_{2},\dots,\gamma_{\lceil n/2\rceil})\ .$ (3)
A design that maximizes WSC is then highly likely to satisfy the two weak sign consistency conditions across many $\mathcal{A}$ where $|\mathcal{A}|\leq\lceil n/2\rceil$. It would be preferred that the left hand sides of (1) and (2) equal 0, again corresponding to orthogonal designs, but nonorthogonal designs may still have a large WSC. The authors employ a hybrid elitist Pareto optimality-based coordinate exchange algorithm (Cao et al., 2017) to identify designs on the Pareto front of the EC and WSC. We refer to their optimal designs as Dantzig-based Consistency Designs (DCDs).
The DCD criteria require evaluations across all $\mathcal{A}$ of sizes $k=2,\dots,\lceil n/2\rceil$. To ease the computational burden, Singh and Stufken (2023) randomly sample subsets for each $k$. Furthermore, their examples initialized their search algorithms with only two designs each, a $\text{Var}(s+)$-optimal design and either a random, $\text{E}(s^{2})$-optimal, or $\text{UE}(s^{2})$-optimal design. The resulting DCDs initialized at a $\text{Var}(s+)$ design were shown via simulation to have better support recovery properties than the $\text{Var}(s+)$-optimal design assuming positive signs of the active effects. However, their improvement over the $\text{Var}(s+)$-optimal design may be attributed to their initialization at that design. That is, they show that the initialized $\text{Var}(s+)$-optimal design is not globally optimal for known signs, but there is no guarantee a DCD is either. Indeed, an orthogonal design is the ideal design under their criteria. It may be that the globally optimal DCD performs worse than the $\text{Var}(s+)$ design. Still, their results provide strong support for constructing SSDs under criteria directly tied to statistical properties of penalized estimators.
Others have considered the role of X in the statistical properties of the lasso, though not in the context of SSDs. Zhao and Yu (2006) identified the strong irrepresentable condition (SIC) on X for establishing support recovery consistency of the lasso (Zhang and Huang, 2008; Jia and Rohe, 2015). Wainwright (2009) and Hastie et al. (2019) have studied support recovery for random designs where the rows of X are independently generated from a zero-mean multivariate Normal distribution. Random designs, however, generally perform poorly for small run designs like SSDs. Based on the lasso’s SIC, Deng et al. (2013) constructed SSDs from nearly orthogonal Latin hypercube designs to minimize the off-diagonals of S assuming factors have settings between $[0,1]$. This is essentially a construction technique for minimizing $\text{UE}(s^{2})$ with more general factor settings, even though SSDs commonly assume fixed settings of ±$\pm 1$. Xing (2015) proposed a lasso SIC criterion to construct optimal two-level SSDs, similar to the $\gamma_{k}$ from (3). Huang et al. (2020) proposed a lasso optimal design theory that applies variance-based criteria to the approximate covariance matrix of the debiased lasso (Javanmard and Montanari, 2014), which is capable of performing inference via confidence intervals. Under the framework of approximate designs (i.e., $n\to\infty$) they note their criteria are not convex and give an equivalence theorem for establishing whether a design is locally optimal. They then propose an algorithmic construction for generating many local optimal approximate designs and implement a rounding procedure on the approximate design’s replication weights to produce an exact design. Overall, their approach requires many stages of approximation, leading to a discrepancy between the approximate and exact design’s covariance matrix.
1.2 Contributions and Overview
This paper addresses the two SSD open problems by developing optimality criteria that directly maximize the probability of sign recovery for the lasso estimator.
The first major contribution of this paper is a local optimal design approach to maximize the sign recovery probability assuming $\boldsymbol{\beta}$ is known and the lasso tuning parameter, $\lambda>0$, is fixed. We then develop criteria that relax the model assumptions and summarize the sign recovery probability across a range of $\lambda$. As optimization of the proposed criteria is challenging, the second major contribution is a criterion based on the $p\times p$ correlation matrix of X. We provide evidence that the ideal correlation matrices under different heuristic criteria are optimal under the new criteria, thus making substantial progress on both open problems. In particular, we prove the surprising result that an orthogonal correlation matrix is suboptimal when $\boldsymbol{z}$ is known. The final contribution is a computationally efficient construction algorithm that leverages the speed of exchange algorithms for optimizing heuristic criteria.
The paper is organized as follows. Section 2 develops exact design optimality criteria targeting the lasso’s sign recovery probability under different assumptions about the model. Section 3 presents a pseudo-approximate optimal design framework that targets the optimal correlation matrix of X and justifies the ideal designs sought after by heuristic criteria under different assumptions about $\boldsymbol{z}$. Section 4 describes a computationally efficient algorithm for constructing exact optimal designs under the proposed criteria. We then demonstrate the optimal SSDs under the new framework in Section 5. We conclude the paper with a discussion in Section 6, describing important implications of the results and future work.
2 Exact Local Optimality Criteria
The lasso is often applied to the centered and scaled design matrix $\textbf{F}=(\textbf{I}-\textbf{P}_{1})\textbf{X}\textbf{V}^{-1/2}$ where $\textbf{P}_{1}=n^{-1}\boldsymbol{1}\boldsymbol{1}^{T}$ and V is a diagonal matrix comprised of the diagonal elements of $n^{-1}\textbf{X}^{T}(\textbf{I}-\textbf{P}_{1})\textbf{X}$. The analysis then targets support/sign recovery of $\boldsymbol{\beta}^{*}=\textbf{V}^{1/2}\boldsymbol{\beta}$. The diagonal elements of $\textbf{V}^{1/2}$ are nonnegative and bounded above by 1, making $|\beta^{*}_{j}|\leq|\beta_{j}|$, and $\text{sign}(\beta_{j}^{*})=\text{sign}(\beta_{j})$ when V has all positive diagonal elements. The support, $\mathcal{A}$, is estimated by the support of the lasso estimator
$\hat{\boldsymbol{\beta}}^{*}=\operatorname*{arg\,min}_{\boldsymbol{\beta}^{*}}\frac{1}{2}\boldsymbol{\beta}^{*T}\textbf{C}\boldsymbol{\beta}^{*}-\frac{1}{n}\boldsymbol{y}^{T}\textbf{F}\boldsymbol{\beta}^{*}+\lambda\sum_{j=1}^{p}|\beta_{j}^{*}|\ ,$
where $\textbf{C}=n^{-1}\textbf{F}^{T}\textbf{F}$ is the correlation matrix of the columns of X. Denote the estimated support and sign of the resulting lasso estimate, $\hat{\boldsymbol{\beta}}^{*}$, by $\hat{\mathcal{A}}=\{j:|\hat{\beta}_{j}^{*}|>0\}$ and $\hat{\boldsymbol{\boldsymbol{z}}}=\text{sign}(\hat{\boldsymbol{\beta}}^{*})$, respectively.
We denote submatrices of X and F corresponding to column subsets $\mathcal{T}\subseteq\{1,\dots,p\}$ by $\textbf{X}_{\mathcal{T}}$ and $\textbf{F}_{\mathcal{T}}$, respectively. For a $p\times 1$ vector, $\boldsymbol{v}$, $\boldsymbol{v}_{\mathcal{T}}$ denotes the $|\mathcal{T}|\times 1$ vector with the $\mathcal{T}$ elements of $\boldsymbol{v}$. For all other matrices, we will consider submatrices by selecting both rows and columns with two index sets $\mathcal{U}$ and $\mathcal{T}$. That is, for a matrix M, let $\textbf{M}_{\mathcal{U}\mathcal{T}}$ denote the submatrix of M with rows and columns indexed by $\mathcal{U}$ and $\mathcal{T}$, respectively. For brevity, we will denote $\textbf{M}_{\mathcal{T}\mathcal{T}}=\textbf{M}_{\mathcal{T}}$, which should not be confused with a subsetting of columns alone.
For known $\hat{\mathcal{A}}$ and $\hat{\boldsymbol{z}}$, the lasso KKT conditions are (Zhao and Yu, 2006; Tibshirani, 2012):
$\hat{\boldsymbol{\beta}}_{\hat{\mathcal{A}}}^{*}$ $=\frac{1}{n}\textbf{C}_{\hat{\mathcal{A}}}^{-1}\textbf{F}_{\hat{\mathcal{A}}}^{T}\boldsymbol{y}-\lambda\textbf{C}_{\hat{\mathcal{A}}}^{-1}\hat{\boldsymbol{z}}_{\hat{\mathcal{A}}}\ ,$ (4)
$\boldsymbol{0}$ $<\hat{\textbf{Z}}_{\hat{\mathcal{A}}}\hat{\boldsymbol{\beta}}_{\hat{\mathcal{A}}}^{*}\ ,$ (5)
$\ \lambda\boldsymbol{1}$ $\geq\left|\textbf{C}_{\hat{\mathcal{I}}\hat{\mathcal{A}}}\hat{\boldsymbol{\beta}}_{\hat{\mathcal{A}}}^{*}-\frac{1}{n}\textbf{F}_{\hat{\mathcal{I}}}^{T}\boldsymbol{y}\right|\ ,$ (6)
where $\textbf{C}_{\hat{\mathcal{A}}}^{-1}=n(\textbf{F}_{\hat{\mathcal{A}}}^{T}\textbf{F}_{\hat{\mathcal{A}}})^{-1}$, $\hat{\textbf{Z}}_{\hat{\mathcal{A}}}=\text{Diag}(\hat{\boldsymbol{z}}_{\hat{\mathcal{A}}})$, and ^=$\hat{\mathcal{I}}=\{j:\hat{\beta}_{j}^{*}=0\}$. For given X, $\boldsymbol{y}$, and $\boldsymbol{\beta}$, we can check whether the resulting $\hat{\boldsymbol{\beta}}^{*}$ has the true support and sign vector by setting $\hat{\mathcal{A}}=\mathcal{A}$ and $\hat{\boldsymbol{z}}_{\hat{\mathcal{A}}}=\boldsymbol{z}_{\mathcal{A}}$ and checking the KKT conditions. If they hold, $\hat{\boldsymbol{\beta}}^{*}$ recovers the sign, which is more stringent than recovering the support.
The proposed criteria in this section rank designs according to the probability of sign recovery, denoted $P(\hat{\boldsymbol{z}}=\boldsymbol{z}\,|\,\textbf{X},\,\boldsymbol{\beta},\,\lambda)$. To calculate the probability, one must specify a $\lambda>0$ and a $\boldsymbol{\beta}$, so the proposed framework resembles optimality frameworks employed for nonlinear models and maximum likelihood estimation (Silvey, 1980; Khuri et al., 2006; Yang and Stufken, 2009). Note that specifying $\boldsymbol{\beta}$ also leads to specifying $\mathcal{A}$, $\mathcal{I}$, and $\boldsymbol{z}$. The criterion $P(\hat{\boldsymbol{z}}=\boldsymbol{z}\,|\,\textbf{X},\,\boldsymbol{\beta},\lambda)$ is a joint probability of two events corresponding to (5) and (6) assuming $\hat{\mathcal{A}}=\mathcal{A}$ and $\hat{\boldsymbol{z}}_{\hat{\mathcal{A}}}=\boldsymbol{z}_{\mathcal{A}}$. The first event corresponds to (5) and checks whether $\text{sign}(\hat{\boldsymbol{\beta}}^{*}_{\mathcal{A}})=\boldsymbol{z}_{\mathcal{A}}$ where $\hat{\boldsymbol{\beta}}^{*}_{\mathcal{A}}$ is from (4) assuming $\hat{\mathcal{A}}=\mathcal{A}$ and $\hat{\boldsymbol{z}}_{\hat{\mathcal{A}}}=\boldsymbol{z}_{\mathcal{A}}$:
$S_{\lambda}$$\coloneqq\left\{\boldsymbol{u}<\sqrt{n}\textbf{Z}_{\mathcal{A}}\textbf{V}^{1/2}_{\mathcal{A}}\boldsymbol{\beta}_{\mathcal{A}}\right\}\ ,$ (7)
where $\boldsymbol{u}\sim N(\lambda\sqrt{n}\,\textbf{Z}_{\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{z}_{\mathcal{A}},\textbf{Z}_{\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\textbf{Z}_{\mathcal{A}})$. The second event corresponds to (6), which checks whether all $j\in\mathcal{I}$ have $\hat{\beta}^{*}_{j}=0$:
$I_{\lambda}\coloneqq\left\{|\boldsymbol{v}|\leq\lambda\sqrt{n}\boldsymbol{1}\right\}\ ,$ (8)
where $\boldsymbol{v}\sim N(\lambda\sqrt{n}\textbf{C}_{\mathcal{I}\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{z}_{\mathcal{A}},\textbf{C}_{\mathcal{I}}-\textbf{C}_{\mathcal{I}\mathcal{A}}\textbf{C}_{\mathcal{A}}^{-1}\textbf{C}_{\mathcal{A}\mathcal{I}})$. The probability of event (8) depends on $\boldsymbol{\beta}$ only through $\boldsymbol{z}$. The probabilities of both events depend on X only through its C and V, a fact we exploit in Section 3. For brevity, we will sometimes refer to these events as simply $S_{\lambda}$ and $I_{\lambda}$.
Remark 1 There is a tradeoff between the probabilities of (7) and (8) so both should be considered simultaneously. Deng et al. (2013) and Xing (2015) focus on heuristic criteria based on the lasso’s SIC, which optimizes (8) alone.
Remark 2 The probability of support recovery, $P(\hat{\mathcal{A}}=\mathcal{A}\,|\,\textbf{X},\,\boldsymbol{\beta},\,\lambda)$, requires consideration of all possible $2^{k}$ sign vectors that have the same 00 elements as $\boldsymbol{z}$ but alternative ±$\pm 1$ elements indexed by $\mathcal{A}$. Defining $\mathcal{Z}_{\mathcal{A}}$ as the set of all such sign vectors, we have $P(\hat{\mathcal{A}}=\mathcal{A}\,|\,\textbf{X},\,\boldsymbol{\beta},\,\lambda)=2^{-k}\sum_{\tilde{\boldsymbol{z}}\in\mathcal{Z}_{\mathcal{A}}}P(\hat{\boldsymbol{z}}=\tilde{\boldsymbol{z}}\,|\,\textbf{X},\,\boldsymbol{\beta},\,\lambda)$. Each $P(\hat{\boldsymbol{z}}=\tilde{\boldsymbol{z}}\,|\,\textbf{X},\,\boldsymbol{\beta},\,\lambda)$ may be calculated by replacing $\boldsymbol{z}_{\mathcal{A}}$ and $\boldsymbol{z}$ with $\tilde{\boldsymbol{z}}_{\mathcal{A}}$ and $\tilde{\boldsymbol{z}}$, respectively, in events (7) and (8).
Remark 3 As $|\beta_{j}^{*}|\to\infty$ for all $j\in\mathcal{A}$, the probability of sign recovery and support recovery are equivalent. To see this, note that $\textbf{Z}_{\mathcal{A}}\textbf{V}^{1/2}_{\mathcal{A}}\boldsymbol{\beta}_{\mathcal{A}}=\textbf{V}^{1/2}_{\mathcal{A}}|\boldsymbol{\beta}_{\mathcal{A}}|$ so $P(S_{\lambda}\,|\,\textbf{C}_{\mathcal{A}},\,\textbf{V}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})\to 1$ since the area of integration approaches $\mathbb{R}^{k}$, the support of $\boldsymbol{u}$. For any other $\tilde{\boldsymbol{z}}\in\mathcal{Z}_{\mathcal{A}}$, $\tilde{\textbf{Z}}_{\mathcal{A}}\textbf{V}^{1/2}_{\mathcal{A}}\boldsymbol{\beta}_{\mathcal{A}}$ would have negative elements, causing $P(S_{\lambda}\,|\,\textbf{C}_{\mathcal{A}},\,\textbf{V}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})\to 0$ because the region of integration for the $u_{j}$ corresponding to these negative elements approaches $(-\infty,\lim_{t\to-\infty}t)$, which has zero probability. Hence $P(\hat{\boldsymbol{z}}=\tilde{\boldsymbol{z}}\,|\,\textbf{X},\,\boldsymbol{\beta},\,\lambda)\to 0$ for any $\tilde{\boldsymbol{z}}\neq\boldsymbol{z}$. Finally, $I_{\lambda}$ depends on $\boldsymbol{\beta}_{\mathcal{A}}$ only through $\boldsymbol{z}_{\mathcal{A}}$ so it is unaffected by increasing $|\beta_{j}^{*}|$, hence $P(\hat{\mathcal{A}}=\mathcal{A}\,|\,\textbf{X},\,\boldsymbol{\beta})\to P(\hat{\boldsymbol{z}}=\boldsymbol{z}\,|\,\textbf{X},\,\boldsymbol{\beta})$.
Remark 4 We assume $\textbf{C}_{\mathcal{A}}^{-1}$ always exists, making $\boldsymbol{u}$ nondegenerate, but $\boldsymbol{v}$ will have a degenerate multivariate Normal distribution for SSDs because C cannot be full rank. The probability of this event can be calculated following some linear transformation of $\boldsymbol{v}$.
2.1 Local criteria assuming known $\boldsymbol{\beta}$
For a fixed $\lambda$ and known $\boldsymbol{\beta}$, define the local optimality criterion
$\phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta})=P(\hat{\boldsymbol{z}}=\boldsymbol{z}\,|\,\textbf{X},\,\boldsymbol{\beta})=P(S_{\lambda}\cap I_{\lambda}\,|\,\textbf{C},\,\textbf{V}_{\mathcal{A}},\,\boldsymbol{\beta})\ .$ (9)
A $\phi_{\lambda}$-optimal design is $\textbf{X}^{*}=\text{argmax}_{\textbf{X}}\ \phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta})$. This approach is impractical, particularly due to its perfect knowledge of $\mathcal{A}$, but it is foundational for the more practical criteria that allow for uncertainty about $\mathcal{A}$. The following is a fundamental result about the role of $\boldsymbol{z}$ in $\phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta})$. Its proof and all future proofs may be found in the Supplementary Materials.
Lemma 1.
For a given X, events $S_{\lambda}$ and $I_{\lambda}$ are independent, and $\phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta})=\phi_{\lambda}(\textbf{X}\,|\,-\boldsymbol{\beta})$ for any $\boldsymbol{\beta}$ and its reflection, $-\boldsymbol{\beta}$. Hence an optimal $\textbf{X}^{*}$ for $\phi_{\lambda}$ under $\boldsymbol{\beta}$ is also optimal for $\phi_{\lambda}$ under $-\boldsymbol{\beta}$.
Lemma 1 includes a design equivalence result between $\boldsymbol{\beta}$ and its reflection, $-\boldsymbol{\beta}$. The following equivalence theorem is more general and further simplifies the implementation of the framework:
Theorem 1.
For a $\boldsymbol{\beta}$ where $\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1}$, consider $\tilde{\boldsymbol{\beta}}=\tilde{\textbf{Z}}\boldsymbol{\beta}$ where $\tilde{\textbf{Z}}$ is any diagonal matrix comprised of ±$\pm 1$. Then for any X, $\phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta})=\phi_{\lambda}(\textbf{X}\tilde{\textbf{Z}}\,|\,\tilde{\boldsymbol{\beta}})$. Consequently, if $\textbf{X}^{*}$ is optimal for $\phi_{\lambda}$ under a $\boldsymbol{\beta}$, then $\tilde{\textbf{X}^{*}}=\textbf{X}^{*}\tilde{\textbf{Z}}$ is optimal for $\phi_{\lambda}$ under $\tilde{\boldsymbol{\beta}}.$
A consequence of Theorem 1 is that, when discussing criteria involving a known sign vector, we can assume $\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1}$ without loss of generality.
The following criterion assumes the magnitudes of the $\beta_{j}$ are known and treats $\boldsymbol{z}_{\mathcal{A}}$ as an unknown quantity and calculates the expected probability of sign recovery assuming all $\tilde{\boldsymbol{z}}\in\mathcal{Z}_{\mathcal{A}}$ are equally likely. Without loss of generality, the magnitudes are expressed by $\boldsymbol{\beta}>\boldsymbol{0}$ and the criterion is defined to be
$\phi_{\lambda}^{\pm}(\textbf{X}\,|\,\boldsymbol{\beta})=\frac{1}{2^{k}}\sum_{\tilde{z}\in\mathcal{Z}_{\mathcal{A}}}\phi_{\lambda}(\textbf{X}\,|\,\tilde{\textbf{Z}}\boldsymbol{\beta})=\frac{1}{2^{k}}\sum_{\tilde{z}\in\mathcal{Z}_{\mathcal{A}}}\phi_{\lambda}(\textbf{X}\tilde{\textbf{Z}}\,|\,\boldsymbol{\beta})\ .$ (10)
Calculating all $2^{k}$ probabilities can be computationally intensive, but Lemma 1 allows us to halve the number of computations. We state this as a corollary:
Corollary 1.
For a fixed $\mathcal{A}$ where $|\mathcal{A}|=k$, let $\mathcal{Z}_{\mathcal{A}}^{\pm}$ denote the subset of $\mathcal{Z}_{\mathcal{A}}$ including all $2^{k-1}$ unique $\boldsymbol{z}$ up to reflection. Then $\phi_{\lambda}^{\pm}(\textbf{X}\,|\,\boldsymbol{\beta})=2^{-(k-1)}\sum_{\tilde{z}\in\mathcal{Z}_{\mathcal{A}}^{\pm}}\phi_{\lambda}(\textbf{X}\tilde{\textbf{Z}}\,|\,\boldsymbol{\beta})$.
We now investigate designs that maximize $\phi_{\lambda}$ and $\phi_{\lambda}^{\pm}$. Knowledge of $\mathcal{A}$ when constructing an optimal design under either criteria leads to a trivial construction for the columns $\textbf{X}_{\mathcal{I}}^{*}$.
Proposition 1.
For a known $\mathcal{A}$, there exists an optimal design $\textbf{X}^{*}$ where $\textbf{X}^{*}_{\mathcal{I}}=\textbf{J}$. Furthermore, the criteria $\phi_{\lambda}(\textbf{X}^{*}\,|\,\boldsymbol{\beta})$ and $\phi_{\lambda}^{\pm}(\textbf{X}^{*}\,|\,\boldsymbol{\beta})$ for such designs depend only on $S_{\lambda}$.
Optimal designs for criteria with perfect knowledge about $\mathcal{A}$ completely confound the columns of the inactive factors with the intercept, making $P(I_{\lambda}\,|\,\textbf{C},\,\boldsymbol{z})=1$. Finding the optimal design, which may originally have been supersaturated, then reduces to finding a full column rank matrix $\textbf{X}_{\mathcal{A}}$ that maximizes probability of the $S_{\lambda}$ event(s).
To demonstrate this new theory, we will now show that if $\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1}$ is assumed known, an orthogonal design may not maximize $P(S_{\lambda}\,|\,\textbf{C}_{\mathcal{A}},\,\textbf{V}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})$. For any $\textbf{X}_{\mathcal{A}}$ the bounds of integration for $S_{\lambda}$ and the mean of $\boldsymbol{u}$ are proportional to $\textbf{V}_{\mathcal{A}}^{1/2}\boldsymbol{\beta}_{\mathcal{A}}$ and $\lambda\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}$, respectively. Hence, a design that maximizes $\textbf{V}_{\mathcal{A}}^{1/2}\boldsymbol{\beta}_{\mathcal{A}}-\lambda\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}$ may maximize $P(S_{\lambda}\,|\,\textbf{C}_{\mathcal{A}},\,\textbf{V}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})$, although the covariance matrix of $\boldsymbol{u}$ also plays a role. For any orthogonal design, this difference is $\boldsymbol{\beta}_{\mathcal{A}}-\lambda\boldsymbol{1}$. We will try to find a $\textbf{X}_{\mathcal{A}}$ with larger $P(S_{\lambda}\,|\,\textbf{C}_{\mathcal{A}},\,\textbf{V}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})$ by identifying a design whose $\textbf{V}_{\mathcal{A}}^{1/2}$ and $\textbf{C}_{\mathcal{A}}$ satisfy
$(\textbf{I}-\textbf{V}_{\mathcal{A}}^{1/2})\boldsymbol{\beta}_{\mathcal{A}}\leq(\textbf{I}-\textbf{C}_{\mathcal{A}}^{-1})\lambda\boldsymbol{1}\ .$ (11)
As $(\textbf{I}-\textbf{V}_{\mathcal{A}}^{1/2})\boldsymbol{\beta}_{\mathcal{A}}\geq\boldsymbol{0}$, we need $(\textbf{I}-\textbf{C}_{\mathcal{A}}^{-1})\lambda\boldsymbol{1}\geq\boldsymbol{0}$. For even $n\geq 6$, construct $\textbf{X}_{\mathcal{A}}$ by choosing $k\leq n-1$ columns from the $n\times n$ matrix
$\left(\begin{array}[]{c|c}2\textbf{I}-\textbf{J}&-\textbf{J}\\ \hline\cr\textbf{J}&\textbf{J}-2\textbf{I}\end{array}\right)\ ,$ (14)
where all I and $\textbf{J}=\boldsymbol{1}\boldsymbol{1}^{T}$ have size $n/2\times n/2$. Then $\textbf{V}_{\mathcal{A}}=(1-4/n^{2})\textbf{I}$. Let $k_{1}$ denote the number of columns chosen from the first $n/2$ columns and $k_{2}=k-k_{1}$ be the number from the remaining columns. If either $k_{1}$ or $k_{2}$ equal 0, $\textbf{C}_{\mathcal{A}}$ will be completely symmetric with positive off-diagonal elements $c=1-4n/(n^{2}-4)$. Thus $(\textbf{I}-\textbf{C}_{\mathcal{A}}^{-1})\lambda\boldsymbol{1}\geq\boldsymbol{0}$ and (11) becomes
$\left(1-\frac{\sqrt{n^{2}-4}}{n}\right)\boldsymbol{\beta}_{\mathcal{A}}\leq\lambda\left(1-\frac{n^{2}-4}{kn^{2}-4k(n+1)+4n}\right)\boldsymbol{1}\ .$
When $k_{1}\geq 1$ and $k_{2}\geq 1$, $\textbf{C}_{\mathcal{A}}$ and $\textbf{C}_{\mathcal{A}}^{-1}$ have a $2\times 2$ block partitioned form with completely symmetric, diagonal block matrices and constant off-diagonal block matrices. Then $\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}=(\xi_{1}\boldsymbol{1}_{k_{1}}^{T},\xi_{2}\boldsymbol{1}_{k_{2}}^{T})^{T}=\boldsymbol{\xi}$, and $\textbf{C}_{\mathcal{A}}\boldsymbol{\xi}=\boldsymbol{1}$, which gives the equations
11 11 $=\rho_{11}\xi_{1}+\rho_{12}\xi_{2}$
11 11 $=\rho_{21}\xi_{1}+\rho_{22}\xi_{2}\ ,$
where $\rho_{ij}>0$ is the unique row sum for the corresponding $k_{i}\times k_{j}$ block matrix of $\textbf{C}_{\mathcal{A}}$. Defining $\tilde{k}_{i}=(n-2k_{i})\geq 0$, we have
$\xi_{1}=\frac{\tilde{k}_{2}(n^{2}-4)}{(n-2)^{2}(k_{1}\tilde{k}_{2}+\tilde{k}_{1}k_{2})+\tilde{k}_{1}\tilde{k}_{2}}\geq 0\ ,$ (15)
with equality if and only if $k_{2}=n/2$. A similar expression holds for $\xi_{2}$ with $\tilde{k}_{1}$ in the numerator of (15), and $\xi_{2}=0$ if and only if $k_{1}=n/2$. Since the diagonal elements of $\textbf{C}_{\mathcal{A}}$ are 1, $\rho_{ii}\geq 1$ and it follows $\boldsymbol{1}-\boldsymbol{\xi}>\boldsymbol{0}$. Hence designs constructed in this way satisfy (11) for some combinations of $\boldsymbol{\beta}_{\mathcal{A}}$ and $\lambda$, but whether this implies higher $P(S_{\lambda}\,|\,\textbf{C}_{\mathcal{A}},\,\textbf{V}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})$ depends on the covariance matrix of $\boldsymbol{u}$.
For example, suppose $n=16$ and we have $p=20$ factors but we know the first $k=8$ factors have nonzero effects and the rest are inactive. We considered three SSDs that set all inactive factors’ columns to $+1$, eliminating any ability to estimate them. The first SSD set the $8$ active factor columns to an orthogonal $\textbf{X}_{\mathcal{A}}$ and the other two SSDs set active effects columns according to (14) with $k_{1}=4$ and $8$, respectively. For all designs, $\textbf{C}_{\mathcal{A}}^{-1}\boldsymbol{1}=\xi\boldsymbol{1}$ where $\xi=1$, $0.1575$, and $0.1607$ for the orthogonal, $k_{1}=4$, and $k_{1}=8$ designs, respectively. For the $k_{1}=4$ and $k_{1}=8$ designs, (11) is equivalent to $\lambda^{-1}\boldsymbol{\beta}_{\mathcal{A}}\leq 53.92\times\boldsymbol{1}$ and $\lambda^{-1}\boldsymbol{\beta}_{\mathcal{A}}\leq 53.72\times\boldsymbol{1}$, respectively. The three designs’ $\phi_{\lambda}$ and $\phi_{\lambda}^{\pm}$ values were compared across a range of $\lambda$ and three scenarios for $\boldsymbol{\beta}_{\mathcal{A}}$: $(0.3,0.4,\dots,1)^{T}$, $\boldsymbol{1}$, and $3\times\boldsymbol{1}$. The results for $\phi_{\lambda}$ are shown in the left panels of Figure 1.
Figure 1: Probability of sign recovery for $n=16$, $p=20$, and $k=8$ for an orthogonal design and two designs constructed from selecting columns from (14) with $k_{1}=4$ and $8$. The left and right panels correspond to $\phi_{\lambda}$ and $\phi_{\lambda}^{\pm}$, respectively.
For all scenarios, there is a range of large $\lambda$ values where all designs have $\phi_{\lambda}=0$, a middle range where the two proposed designs outperform the orthogonal design, and a range of small $\lambda$ values where the orthogonal design is superior. The orthogonal design improves over the other two designs before condition (11) is violated, due to the role of the covariance matrices. The improvement is negligible when $\boldsymbol{\beta}_{\mathcal{A}}=3\times\boldsymbol{1}$, implying the covariance matrix is less important as the elements of $\boldsymbol{\beta}_{\mathcal{A}}$ increases. The results for $\phi_{\lambda}^{\pm}$ are shown in the right panels of Figure 1, and clearly favor the orthogonal design.
2.2 Relaxed local criteria
To make the criteria in Section 2.1 more practical, we introduce relaxations regarding the choice of $\lambda$ and the assumptions of the underlying model. To reduce the dependency on $\lambda$, we propose two summary measures of $\phi_{\lambda}$ and $\phi_{\lambda}^{\pm}$. The first summary takes the maximum sign recovery probability: $\phi_{\text{max}}(\textbf{X}\,|\,\boldsymbol{\beta})=\max_{\lambda>0}\phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta})$. However, in practice one will perform some tuning parameter selection strategy to choose a $\lambda$, and so the benefits of a design that optimizes $\phi_{\text{max}}$ will only be realized if the strategy reliably picks the corresponding $\lambda$ that maximizes $\phi_{\lambda}$. Hence, we would like a design that maintains large probabilities for a wide range of $\lambda$. It is common to calculate the lasso solution path with respect to $\log(\lambda)$. Such a transformation stretches the region of $\lambda\in(0,1)$, whose corresponding lasso estimates would receive the least amount of penalization. Therefore, we propose the criterion
[…]
The definitions for $\phi_{\max}^{\pm}(\textbf{X}\,|\,\boldsymbol{\beta})$ and $\phi_{\Lambda}^{\pm}(\textbf{X}\,|\,\boldsymbol{\beta})$ relax assumptions about $\boldsymbol{z}_{\mathcal{A}}$ and measure the maximum expected sign recovery probability and the integrated expected sign recovery probability, assuming all $\boldsymbol{z}_{\mathcal{A}}$ are equally likely.
Turning to relaxing assumptions about the $\boldsymbol{\beta}$, we first assume $|\boldsymbol{\beta}_{\mathcal{A}}|\geq\beta\boldsymbol{1}$ where $\beta$ is the minimum practically significant effect size. Such a $\beta$ value is commonly chosen while performing power calculations so we will calculate the support recovery probability under the most challenging scenario of $|\boldsymbol{\beta}_{\mathcal{A}}|=\beta\boldsymbol{1}$. We will defer the discussion about incorporating uncertainty about $|\boldsymbol{\beta}_{\mathcal{A}}|$ to Section 6. Finally, to address uncertainty about $\mathcal{A}$, we fix $k\leq n-1$ and assume all supports of this size are equally likely. Denote this set of supports by $\mathcal{A}_{k}$.
With the above relaxed assumptions, we first introduce a criterion that incorporates uncertainty about $\mathcal{A}$, but treats $\boldsymbol{z}_{\mathcal{A}}$ as known. This resembles scenarios where $\text{Var}(s+)$-optimal designs have been shown to outperform $\text{E}(s^{2})-$ and $\text{UE}(s^{2})-$optimal SSDs in simulation studies. Assume $z_{j}$ depends only on whether $j\in\mathcal{A}$ or $\mathcal{I}$ as follows. Let $\boldsymbol{z}^{*}$ be the $p\times 1$ vector comprised of ±$\pm 1$ for each $j$, assuming $j\in\mathcal{A}$. Then $z_{j}=z_{j}^{*}$ when $j\in\mathcal{A}$ and $z_{j}=0$ when $j\in\mathcal{I}$. Following Theorem 1, we assume without loss of generality that $\boldsymbol{z}^{*}=\boldsymbol{1}$ to identify the optimal design and then multiply the columns of the design by their actual corresponding $z_{j}^{*}$. For a given $\mathcal{A}$, let A be the $p\times p$ diagonal matrix of all zeroes except for the diagonal elements corresponding to $\mathcal{A}$, which are set to 1. Then the true sign vector is $\boldsymbol{z}=\textbf{A}\boldsymbol{1}$ and the sign-dependent criterion for a given $\beta$ is
$\Phi_{\lambda}(\textbf{X}\,|\,k,\beta)=\binom{p}{k}^{-1}\sum_{\mathcal{A}\in\mathcal{A}_{k}}\phi_{\lambda}(\textbf{X}\,|\,\boldsymbol{\beta}=\beta\textbf{A}\boldsymbol{1})\ ,$
being the expected probability across $\mathcal{A}_{k}$ for a fixed $\lambda$. The notation $\Phi$ is intended to reflect consideration of all $\mathcal{A}\in\mathcal{A}_{k}$. The sign-independent criterion, $\Phi_{\lambda}^{\pm}(\textbf{X}\,|\,k,\beta)$, is similarly defined, considering all $2^{k}$ sign vectors as equally likely. The two summary measures originally defined on $\phi_{\lambda}$ and $\phi_{\lambda}^{\pm}$ are straightforward to define on these new criteria and share similar notation, i.e., $\Phi_{\text{max}}(\textbf{X}\,|\,k,\beta)$ and $\Phi_{\Lambda}(\textbf{X}\,|\,k,\beta)$. Computational details of the criteria may be found in the Supplementary Materials.
Optimizing any of these sign recovery criteria is both analytically and algorithmically challenging. Indeed, simply evaluating such criteria for a given X can be cumbersome. We next discuss properties of these criteria and, motivated by approximate design theory (Pukelsheim, 2006), we propose a more tractable framework centered on finding optimal C matrices.
3 Approximate Local Optimality Criteria
Popular design criteria for least-squares estimation are functions of the eigenvalues of $\textbf{M}=\textbf{X}^{T}(\textbf{I}-\textbf{P}_{1})\textbf{X}$. Such criteria are difficult to optimize for fixed $n$ because of the discrete nature of the problem. A more tractable approach optimizes the criteria but with an extended domain of all symmetric, positive definite matrices, denoted $\mathcal{M}$. Such “approximate” criteria are log concave with respect to $\mathcal{M}$, and are invariant to simultaneous row/column permutations and sign transformations of M. These properties can be exploited to find optimal forms of M. For an $\textbf{M}\in\mathcal{M}$, define its matrix-averaged form to be $\overline{\textbf{M}}=(p!)^{-1}\sum_{\boldsymbol{\Pi}\in\mathcal{P}}\boldsymbol{\Pi}\,\textbf{M}\,\boldsymbol{\Pi}^{T}$ where $\mathcal{P}$ is the set of all $p\times p$ permutation matrices. Then $\overline{\textbf{M}}\in\mathcal{M}$ is a completely symmetric matrix, having constant diagonal elements and constant off-diagonal elements. Further averaging of $\overline{\textbf{M}}$ across all $2^{p}$ sign transformations leads to a matrix proportional to the identity matrix. For a log concave criterion, the criterion value for $\overline{\textbf{M}}$ is greater than or equal to that for M. Hence the search for the optimum M may be restricted to the class of completely symmetric matrices in $\mathcal{M}$. In this class, an optimal $M^{*}$ can be identified, and one can search among exact designs to find ones whose M’s are close to $\textbf{M}^{*}$ (Pukelsheim and Rieder, 1992).
[…]
To simplify our investigation of these new approximate lasso criteria, we condition optimization across the space of positive definite correlation matrices. Since such matrices cannot occur with an SSD, a more appropriate framework tailored for SSDs would require optimization across rank-deficient correlation matrices, which to our knowledge has received little attention in the statistical literature. However, similar to the motivation behind the $\text{E}(s^{2})$- and $\text{UE}(s^{2})$-criterion, an optimal positive definite correlation matrix still provides a hypothetical target which we use in Section 4 to efficiently construct new SSDs to maximize lasso sign recovery probabilities.
The primary goal of this section is to identify optimal, positive definite correlation matrices under the relaxed local criteria. Such criteria are permutation invariant due to $|\boldsymbol{\beta}_{\mathcal{A}}|=\beta\boldsymbol{1}$, which suggests there exists an optimal $\textbf{C}^{*}$ that is completely symmetric. For $p\geq k$, let $\mathcal{C}_{p,k}=\{\textbf{C}=(1-c)\textbf{I}_{p}+c\textbf{J}_{p}\,|\,-(k-1)^{-1}<c<1\}$ denote a set of $p\times p$ completely symmetric correlation matrices. This space does include rank-deficient correlation matrices, corresponding to $-(k-1)^{-1}<c\leq-(p-1)^{-1}$. For $k=p$, we will denote this set as simply $\mathcal{C}_{p}$. If the approximate criteria were log concave across the space of all positive definite correlation matrices, then, similar to eigenvalue-based criteria, matrix averaging implies the optimal correlation matrix must lie in $\mathcal{C}_{p}$. This reduces the dimension of the optimization problem from $\binom{p}{2}$ off-diagonal elements of C to a single off-diagonal value, $c$. Investigating log concavity for approximate criteria is challenging because C is involved in both the covariances and means for $S_{\lambda}$ and $I_{\lambda}$. Derivatives are intractable even for a fixed $\lambda$.
To simplify the investigation of log concavity, Section 3.1 assumes a known $\boldsymbol{\beta}$ and the optimal design has $\textbf{X}_{\mathcal{I}}=\textbf{J}$. This makes the criterion equal to $P(S_{\lambda}\,|\,\textbf{C}_{\mathcal{A}},\,\textbf{V}_{\mathcal{A}},\,\boldsymbol{\beta}_{\mathcal{A}})$ and so we explore the log concavity of this probability with respect to $\textbf{C}_{\mathcal{A}}$. We identify situations when such functions fail to be log concave, but provide evidence that the probability may still be optimized by $\textbf{C}_{\mathcal{A}}\in\mathcal{C}_{k}$. These numerical results motivate focusing optimization of relaxed criteria across $\mathcal{C}_{p,k}$ in Section 3.2.
3.1 Optimal $\textbf{C}_{\mathcal{A}}$ for known $\boldsymbol{\beta}$
Following Proposition 1, approximate local criteria for a known $\boldsymbol{\beta}$ should assume $\textbf{X}_{\mathcal{I}}=\textbf{J}$ and so should be defined only on the space of positive definite $k\times k$ correlation matrices, i.e., $\phi_{\lambda}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})=P(S_{\lambda}\,|\,\textbf{C}_{\mathcal{A}},\,\textbf{I},\,\boldsymbol{\beta}_{\mathcal{A}})$ and $\phi_{\lambda}^{\pm}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})=2^{-(k-1)}\sum_{\tilde{z}\in\mathcal{Z}^{\pm}_{\mathcal{A}}}P(S_{\lambda}\,|\,\tilde{\textbf{Z}}\textbf{C}_{\mathcal{A}}\tilde{\textbf{Z}},\,\textbf{I},\,|\boldsymbol{\beta}_{\mathcal{A}}|)$. To focus on relaxed criteria, we simplified our investigation of log concavity to the case of $\boldsymbol{\beta}_{\mathcal{A}}=\beta\boldsymbol{1}_{k}$. While such criteria are impractical by themselves, they contribute to understanding of more useful relaxed local criteria. The purpose of this section is to investigate when such criteria are log concave and, if they are not, when they may be optimized by a $\textbf{C}_{\mathcal{A}}\in\mathcal{C}_{k}$. We focus on $k=2,3,$ and $8$, and show that $\phi_{\lambda}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})$ is generally optimized by a $\textbf{C}_{\mathcal{A}}\in\mathcal{C}_{k}$, even when it fails to be log concave (which tends to only happen as $\beta$ and $\lambda$ approach 0). The criteria $\phi_{\lambda}^{\pm}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})$ often fail to be log concave for larger values of $\lambda$ but still we are able to find a $\textbf{C}_{\mathcal{A}}\in\mathcal{C}_{k}$ that maximizes (or nearly maximizes) $\phi_{\lambda}^{\pm}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})$.
3.1.1 Examples with $k=2$
With $k=2$ we have an optimization problem of a single correlation value, $|c|<1$, so $\textbf{C}_{\mathcal{A}}^{*}$ must be completely symmetric. Establishing log concavity in this case is essentially investigating the uniqueness of the optimal $c^{*}$, although log concavity is only a sufficient condition for uniqueness. In the Supplementary Materials, we show
$\phi_{\lambda}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})$ $=[2G(\mu(c))-1]\times G((1+c)\beta-\lambda)+2T\left((1+c)\beta-\lambda,\sqrt{\frac{1-c}{1+c}}\right)\ ,$ (16)
where $\mu(c)=\sqrt{(1-c)/(1+c)}((1+c)\beta-\lambda)$; $g(\cdot)$ and $G(\cdot)$ refer to the standard Normal probability density function and cumulative distribution function, respectively; and $T(\cdot,\cdot)$ is Owen’s $T$ function. There is no closed form expression for $T(\cdot,\cdot)$ except in special cases. One such case results from taking the limit:
$\lim_{\beta,\lambda\to 0}\phi_{\lambda}(\textbf{C}\,|\,\boldsymbol{\beta})=2T\left(0,\sqrt{\frac{1-c}{1+c}}\right)=\frac{1}{\pi}\arctan\left(\sqrt{\frac{1-c}{1+c}}\right)\ .$ (17)
The log of (17) is shown in Figure 2 which is decreasing in $c$ and convex for $c<-0.4422$. This shows that $P(S_{\lambda}\,|\,\textbf{C}_{\mathcal{A}},\,\textbf{I},\,\boldsymbol{\beta}_{\mathcal{A}})$ is not log concave for all $\boldsymbol{\beta}_{\mathcal{A}}$ and $\lambda$, although the optimal $c$ is unique.
Figure 2: Plot of the logarithm of equation (17). The function fails to be concave for $c<-0.4422$.
The limiting case in (17) is uninteresting because practical screening problems typically assume $\beta\geq 1$. Moreover, optimal designs under relaxed criteria will not set $\textbf{X}_{\mathcal{I}}=\textbf{J}$, and for such designs, $P(I_{\lambda}\,|\,\textbf{C},\boldsymbol{z})$ will be close to 0 as $\lambda\to 0$, thereby eliminating the benefits of setting $c$ close to $-1$. Finally, for $\beta$ close to 0, sign recovery probability will be significantly smaller than the support recovery probability, and the latter should be preferred for comparing designs.
A general assessment of log concavity may be accomplished by analyzing the second-order derivatives of $\log[\phi_{\lambda}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})]$. However, the derivatives are tedious and too complicated for establishing concavity. Another technique to establish log concavity of a function involving an integral is based on the Prékopa–Leindler inequality (Prékopa, 1971) which requires the integrand to be bivariate log concave. An example is provided in the Supplementary Materials to show this sufficient condition does not hold for the integrands involved in $P(S_{\lambda}\,|\,\textbf{C}_{\mathcal{A}},\,\textbf{I},\,\boldsymbol{\beta}_{\mathcal{A}}=\beta\boldsymbol{1})$. Instead, we numerically investigated the log concavity conjecture under various combinations of $\boldsymbol{\beta}_{\mathcal{A}}$ and $\lambda$. We constructed a maximin design with 750 points to fill the space of possible $\beta$ and $\log(\lambda)$, where $\beta\in[0.001,3]$ and $\log(\lambda)\in[-10,1.2]$. For each setting, we estimated the second-order derivatives of $\log(\phi_{\lambda}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}}))$ for a dense grid of $c\in[-0.99,0.99]$ using second-order finite differences with a step size of $h=0.01$. If the largest value was greater than 0 the corresponding criterion cannot be log concave. Of the 750 settings considered, 49 were not log concave. These cases resembled the limiting case from (17), having $\lambda<0.01$ and $\beta\leq 0.3720$. The optimal $c$ value in all of these cases was unique and close to $c^{*}=-0.99$. Two examples are shown in Figure 3.
Figure 3: Plot of the logarithm of equation (16) for two of the 750 maximin scenarios that violate log concavity. The red line corresponds to the $c^{*}$ value, being $-0.99$ and $-0.94$ for the left and right panels, respectively.
In the Supplementary Materials, we show $\log[\phi_{\lambda}^{\pm}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})]$ is symmetric about $c=0$. For it to be log concave, it is necessary (but not sufficient) for $c^{*}=0$. Of the 750 maximin settings, 81 failed to be log concave across a range of $\beta\in(0.01,3)$ and $\log(\lambda)\in(-9.854,1.198)$. Two examples that violate log concavity are shown in Figure 4, with one case having $c^{*}=0$. The lack of log concavity is not surprising, as the criterion is an average of functions, which can fail to be log concave even when the individual functions themselves are log concave. If log concavity were a necessary property, one could replace the average with the minimum probability across all sign vectors. We provide some discussion on this alternative definition in Section 6.
Figure 4: Plot of $\log[\phi_{\lambda}^{\pm}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})]$ for two of the 750 maximin scenarios that violate log concavity. The red line corresponds to the $c^{*}$ value, being 00 and $-0.80$ for the left and right panels, respectively.
3.1.2 Examples with $k=3$
Maximizing $\phi_{\lambda}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})$ when $k=3$ requires optimization across three correlation parameters, $(c_{12},c_{13},c_{23})$. Such criteria are permutation invariant and so, if they are log concave, then $\log[\phi_{\lambda}(\textbf{C}_{\mathcal{A}}\,|\,\beta_{\mathcal{A}})]\leq\log[\phi_{\lambda}(\bar{\textbf{C}}_{\mathcal{A}}\,|\,\beta_{\mathcal{A}})]$. If there exists a $\textbf{C}_{\mathcal{A}}$ where this does not hold, then $\phi_{\lambda}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})$ cannot be log concave. For each of the 750 settings from Section 3.1.1, we randomly generated 500 full-rank correlation matrices and checked how many violated $\log[\phi_{\lambda}(\textbf{C}_{\mathcal{A}}\,|\,\beta_{\mathcal{A}})]\leq\log[\phi_{\lambda}(\bar{\textbf{C}}_{\mathcal{A}}\,|\,\beta_{\mathcal{A}})]$. We also recorded the best observed $\textbf{C}_{\mathcal{A}}$ matrix and compared its criterion value to $\textbf{C}_{\mathcal{A}}^{*}\in\mathcal{C}_{3}$, the optimal correlation matrix in $\mathcal{C}_{3}$.
Of the 750 settings considered, 87 found at least one $\textbf{C}_{\mathcal{A}}$ that violated $\log[\phi_{\lambda}(\textbf{C}_{\mathcal{A}}\,|\,\beta_{\mathcal{A}})]\leq\log[\phi_{\lambda}(\bar{\textbf{C}}_{\mathcal{A}}\,|\,\beta_{\mathcal{A}})]$. All such settings had a $\beta<0.628$ and $\lambda<0.028$, similar to the scenarios that violated log concavity for $k=2$. While these 87 settings were not log concave, all of the maximum observed $\textbf{C}_{\mathcal{A}}$ were improved upon by a $\textbf{C}_{\mathcal{A}}\in\mathcal{C}_{3}$ with off-diagonal $c^{*}$ nearly equal to the lower bound of $-0.5$. Therefore, even in the observed cases where log concavity does not hold, it is reasonable to restrict the search for an optimal correlation matrix to $\mathcal{C}_{3}$.
Given what we observed for the $k=2$ case, we did not anticipate $\phi_{\lambda}^{\pm}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})$ to be log concave even for completely symmetric C, except for certain combinations of $\beta$ and $\lambda$. Our main concern though is whether there exists an optimal $\textbf{C}_{\mathcal{A}}$ that is completely symmetric. We performed the same numerical study as for the known sign case but checked whether $\log[\phi_{\lambda}^{\pm}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})]\leq\log[\phi_{\lambda}^{\pm}(\bar{\textbf{C}}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})]$ and, if not, whether the best observed $\textbf{C}_{\mathcal{A}}$ could be improved upon by some $\textbf{C}_{\mathcal{A}}^{*}\in\mathcal{C}_{3}$. Only 65 of the 750 settings violated the inequality with all but one case having $\lambda>1.3$. Of these 65 settings, $\textbf{C}_{\mathcal{A}}^{*}\in\mathcal{C}_{3}$ improved the criterion value over the best observed $\textbf{C}_{\mathcal{A}}$ in all but one case. This one case had $\beta=2.8651$ and $\lambda=2.7535$ and the optimal observed off-diagonals were $(c_{12},c_{13},c_{23})=(-0.0053,-0.0484,-0.5832)$ with $\phi_{\lambda}^{\pm}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})=0.1729$; the criterion value for $\textbf{C}_{\mathcal{A}}^{*}\in\mathcal{C}_{3}$ was $0.1686$. This numerical study provides evidence that for a fixed $\lambda$, the optimal correlation matrix may not be completely symmetric, but generally there exists such a matrix that can perform nearly as well or better.
We also evaluated log concavity of $\phi_{\lambda}^{\pm}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})$ across the 750 settings restricted to $\textbf{C}_{\mathcal{A}}\in\mathcal{C}_{3}$, calculating the second-order derivatives of $\log[\phi_{\lambda}^{\pm}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})]$ with second-order finite differences with a step size of $h=0.01$. Of the 750 settings, we found 77 that violated the log concavity condition but all scenarios gave a unique $c^{*}$. Unlike with $\phi_{\lambda}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})$, only three of theses 77 settings had $\lambda<0.4$, each with an optimal $c^{*}=0$ that failed to be log concave at $c$ close to 1. The remaining 74 settings had $\lambda>0.8$ and either had $c^{*}\in(0.5,0.8)$ (with a maximum second-order derivative at $c\in(0,0.241)$) or $c^{*}=0$ (with a maximum second-order derivative at $c\in(0.190,0.811))$. When $c^{*}>0$, the maximal probability was essentially driven by only the probability of recovering $\boldsymbol{z}=\boldsymbol{1}$, with all other sign recovery probabilities essentially 00. This is a consequence of averaging over sign vectors rather than focusing on the minimum probability across the sign vectors.
3.1.3 Examples with $k=8$
Maximizing $\phi_{\lambda}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})$ when $k=8$ involves an optimization problem across 28 correlation parameters. We focus our numerical investigation to $\beta=4(=\sqrt{16}\times 1)$ and $\beta=12(=\sqrt{16}\times 3)$, which correspond to the last two panel rows in Figure 1 in Section 2.1. These two values are often encountered in simulation studies performed in SSD literature. For each $\beta$ we checked the log concavity inequality $\log[\phi_{\lambda}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})]\leq\log[\phi_{\lambda}(\bar{\textbf{C}}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})]$ across 1000 randomly generated positive definite correlation matrices for $\log(\lambda)$ between $-3$ to $3$ in steps of $0.2$. We also checked whether the maximum observed probability could be improved upon by $\textbf{C}_{\mathcal{A}}^{*}\in\mathcal{C}_{8}$. Finally, we repeated the numerical investigation for $\phi_{\lambda}^{\pm}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})$ but for only 500 randomly generated positive definite correlation matrices due to the increased computations from averaging over $2^{7}$ sign vectors.
For $\beta=4$, the inequality constraint was violated for 29 out of the 31 possible $\lambda$ values. For $\log(\lambda)\leq 0.2$, no $\textbf{C}_{\mathcal{A}}\in\mathcal{C}_{8}$ had a higher $\phi_{\lambda}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})$ than the maximum observed probability, but the differences in probabilities were negligible, with both close to 1. For $\log(\lambda)>0.2$, we were able to find a $\textbf{C}_{\mathcal{A}}\in\mathcal{C}_{8}$ with a higher $\phi_{\lambda}(\textbf{C}\,|\,\boldsymbol{\beta})$ than the observed maximum, with meaningful differences in the two probabilities. A notable instance was for $\log(\lambda)=2$ where the maximum observed probability was $0.1282$ and the maximum probability across $\mathcal{C}_{8}$ was $0.7141$. Similar behavior was observed for $\beta=12$, except we found a $\textbf{C}_{\mathcal{A}}\in\mathcal{C}_{8}$ that exceeded the maximum observed probability for all $\lambda$.
Our investigation into $\phi_{\lambda}^{\pm}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})$ revealed similar results as for $k=3$, being that the log concavity condition is violated for larger $\lambda$. For $\beta=4$, this occurred for $\log(\lambda)\in\{1.6,1.8,...,2.4\}$ but we were always able to find a $\textbf{C}_{\mathcal{A}}^{*}\in\mathcal{C}_{8}$ that had larger $\phi_{\lambda}^{\pm}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})$. For these cases, the off-diagonal element for the $\textbf{C}_{\mathcal{A}}^{*}$’s were $c^{*}\geq 0.4172$, giving the designs a slight edge in recovering $\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1}$. Their $\phi_{\lambda}^{\pm}(\textbf{C}_{\mathcal{A}}\,|\,\boldsymbol{\beta}_{\mathcal{A}})$ values, however, were still close to 0, due to the fact that $\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1}$ is only one of $128$ possible sign vectors. We observed similar results for $\beta=12$ for $\log(\lambda)=\{2.6,2.8,3\}$.
3.2 Optimal $\textbf{C}\in\mathcal{C}_{p,k}$ for relaxed criteria
The numerical results in Section 3.1 provide a reasonable justification that the optimal $\textbf{C}^{*}$ (or at least a nearly optimal $\textbf{C}^{*}$) for approximate relaxed criteria can be found in $\mathcal{C}_{p,k}$. Therefore, this section focuses on identifying such a $\textbf{C}^{*}$. This optimization can be done numerically and we investigate the localized derivatives of the criteria at $c=0$ for fixed $\lambda$ values. We argue $\textbf{C}^{*}=\textbf{I}$ for summaries of $\Phi_{\lambda}^{\pm}(\textbf{C}\,|\,k,\beta)$, and that there exists a $\textbf{C}^{*}$ with all positive off-diagonals for summaries of the sign-dependent $\Phi_{\lambda}(\textbf{C}\,|\,k,\beta)$.
Events (7) and (8) will be denoted here by $P(S_{\lambda}\,|\,c,\boldsymbol{\beta}_{\mathcal{A}})$ and $P(I_{\lambda}\,|\,c,\boldsymbol{z})$. The probabilities for the two events for $\textbf{C}\in\mathcal{C}_{p,k}$ are given in the following lemma:
Lemma 2.
[…]
$\boldsymbol{u}$ $\sim N\left(\frac{\lambda}{1-c}\left[\boldsymbol{1}-z_{\mathcal{A}}\gamma\boldsymbol{z}_{\mathcal{A}}\right],\ \frac{1}{1-c}\left[\textbf{I}_{k}-\gamma\boldsymbol{z}_{\mathcal{A}}\boldsymbol{z}_{\mathcal{A}}^{T}\right]\right)\ ,$
$\boldsymbol{v}$ $\sim N\left(\lambda z_{\mathcal{A}}\gamma\boldsymbol{1},(1-c)\left[\textbf{I}+\gamma\textbf{J}\right]\right)\ ,$
with $z_{\mathcal{A}}=\boldsymbol{1}^{T}\boldsymbol{z}_{\mathcal{A}}$ and $\gamma=c/(1+c(k-1))$. Moreover, if $\boldsymbol{\beta}_{\mathcal{A}}$ does not depend on $\mathcal{A}$, the probabilities of the two events are constant across all such $\mathcal{A}$.
When $\boldsymbol{\beta}_{\mathcal{A}}=\beta\boldsymbol{1}$ the random vectors $\boldsymbol{u}$ and $\boldsymbol{v}$ are distributed as
$\boldsymbol{u}\sim N\left(\frac{\lambda}{1+c(k-1)}\boldsymbol{1},\frac{1}{1-c}\left[\textbf{I}-\gamma\textbf{J}\right]\right),\qquad\boldsymbol{v}\sim N\left(\lambda k\gamma\boldsymbol{1},(1-c)\left[\textbf{I}+\gamma\textbf{J}\right]\right)\ ,$
for all $\mathcal{A}\in\mathcal{A}_{k}$. The corresponding criterion defined with respect to $\mathcal{C}_{p,k}$ and for a fixed $\lambda$ is denoted
$\psi_{\lambda}(c\,|\,k,\beta)=P(S_{\lambda}\,|\,c,\boldsymbol{\beta}_{\mathcal{A}}=\beta\boldsymbol{1})\times P(I_{\lambda}\,|\,c,\boldsymbol{z})\ .$ (18)
The analog to $\Phi_{\lambda}^{\pm}(\textbf{X}\,|\,k,\beta)$ is denoted
$\psi_{\lambda}^{\pm}(c\,|\,k,\beta)=\frac{1}{2^{k-1}}\sum_{\tilde{\boldsymbol{z}}\in\mathcal{Z}_{\mathcal{A}}^{\pm}}P(S_{\lambda}\,|\,c,\boldsymbol{\beta}_{\mathcal{A}}=\beta\tilde{\boldsymbol{z}})\times P(I_{\lambda}\,|\,c,\boldsymbol{z}=\tilde{\boldsymbol{z}})\ .$ (19)
The summarized versions of (18) and (19) across $\lambda$ will be denoted similarly to their exact design counterparts from Section 2, replacing $\Phi$ with $\psi$.
The criteria $\psi_{\lambda}$ and $\psi_{\lambda}^{\pm}$, as well as their summarized versions, involve a single variable but are still challenging to optimize analytically. Fortunately, numerical optimization of these criteria is straightforward and computationally efficient. We demonstrate numerical optimization of these new criteria for the situation of $p=10$, $k=4$, and $\beta=2\sqrt{10}$. Figure 5 shows the contour plots of $\psi_{\lambda}$ and $\psi_{\lambda}^{\pm}$. The optimal $c$ values for the corresponding integral summary measures, $\psi_{\Lambda}$ and $\psi_{\Lambda}^{\pm}$, are $0.14$ and 00, respectively. The resulting optimal C matrices then match the ideal forms for the $\text{Var}(s+)$ and $\text{UE}(s^{2})$-criterion. While these ideal C are not possible for SSDs, this example provides further justification for these two heuristic SSD criteria for sign recovery under the lasso.
Figure 5: Contour plots of sign recovery probabilities when C is completely symmetric across possible values of the off-diagonal $c$ and $\log(\lambda)$ where $k=4$ factors are active with $\beta=2\sqrt{10}$. The red lines correspond to values of $c=0.14$ for $\psi_{\Lambda}$ and $c=0$ for $\psi_{\Lambda}^{\pm}$.
[…]
We now study the behavior of $\psi_{\lambda}$ and $\psi^{\pm}_{\lambda}$ in a neighborhood about $c=0$ to address the potential optimality of orthogonal designs. Direct application of the multivariate Leibniz integral rule gives the following lemma.
Lemma 3.
For all $\mathcal{A}$ where $|\mathcal{A}|=k$ and $\boldsymbol{\beta}_{\mathcal{A}}=\beta\boldsymbol{1}>\boldsymbol{0}$, $\frac{d}{dc}P(I_{\lambda}|c,\boldsymbol{z})\big{|}_{c=0}=0$ for all $\lambda$ and $\frac{d}{dc}P(S_{\lambda}|c,\boldsymbol{\beta}_{\mathcal{A}}=\beta\boldsymbol{1})\big{|}_{c=0}\geq 0$ for $\lambda$ satisfying
$2\lambda$ $\geq\frac{g(\tau)}{G(\tau)}\ ,$ (20)
where $g(\cdot)$ and $G(\cdot)$ refer to the standard Normal probability density function and cumulative distribution function, respectively, and $\tau=\beta-\lambda$.
A direct consequence of Lemma 3 is that $\psi_{\lambda}(0|k,\beta)$ can often be improved upon by some $c>0$. We state this important result as a theorem.
Theorem 2.
For all $\mathcal{A}$ where $|\mathcal{A}|=k$, $\frac{d}{dc}\psi_{\lambda}(c|k,\beta)\big{|}_{c=0}>0$ for $\lambda$ satisfying (20). Hence there exists some $c_{\lambda}>0$, where $\psi_{\lambda}(c_{\lambda}|k,\beta)>\psi_{\lambda}(0|k,\beta)$.
Applying this result to $\psi_{\lambda}(c\,|\,k,\beta)$ shown in Figure 5, inequality (20) holds for $\log(\lambda)\geq-22.763$, which covers the entire region of $\lambda$ values for which $\psi_{\lambda}(c\,|\,k,\beta)>0$. It follows then that some $c>0$ will maximize the summary measures $\psi_{\Lambda}(c\,|\,k,\beta)$ or $\psi_{\max}(c\,|\,k,\beta)$. Theorem 2 thus provides a mathematically rigorous justification that the ideal designs sought after by the $\text{Var}(s+)$-criterion maximize the probability of sign recovery for $\boldsymbol{\beta}_{\mathcal{A}}=\beta\boldsymbol{1}$. The investigations in Section 3.1 further suggest such designs may be superior when $\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1}$, so long as the elements of $\boldsymbol{\beta}_{\mathcal{A}}$ are sufficiently large.
The following theorem establishes a similar justification for heuristic orthogonality measures under unknown signs.
Theorem 3.
For all $\mathcal{A}$ where $|\mathcal{A}|=k$ and $\beta>0$, $\frac{d}{dc}\psi_{\lambda}^{\pm}(c\,|\,k,\beta)\big{|}_{c=0}=0$ for all $\lambda$. Moreover, $c=0$ is a local maximum for $\psi_{\lambda}^{\pm}(c\,|\,k,\beta)$ when
$\begin{split}\frac{q}{\binom{k}{2}}\ \frac{\lambda g(\lambda)}{1-2G(-\lambda)}&\left(k(1-\lambda^{2})+(q-1)\frac{\lambda g(\lambda)}{1-2G(-\lambda)}\right)\leq\\ &\frac{g(\tau)}{G(\tau)}\left(\beta+\lambda+\lambda^{2}\tau-\left[\frac{\beta^{2}-\lambda^{2}}{2}+\beta\lambda\right]\frac{g(\tau)}{G(\tau)}\right)\end{split}$ (21)
where $q=p-k$.
Applying this result to $\psi_{\lambda}^{\pm}(c\,|\,k,\beta)$ shown in Figure 5, inequality (21) holds for $\log(\lambda)\in[-0.988,0.640]$. For $\log(\lambda)$ outside this region, there are some $c>0$ for which $\psi_{\lambda}^{\pm}(c)>0$, although the probabilities are relatively small. These small probabilities do not influence $\psi_{\max}^{\pm}$ and have minimal influence on $\psi_{\Lambda}^{\pm}$, making $c=0$ a likely global maximum for both criteria. We conjecture that generally $c^{*}=0$ and some $c^{*}>0$ are global maxima for $\psi_{\lambda}^{\pm}(c\,|\,k,\beta)$ and $\psi_{\lambda}(c\,|\,k,\beta)$, respectively.
4 Fast Design Construction with HILS
Section 3 identified optimal forms of completely symmetric C matrices under different assumptions about $\boldsymbol{z}$. Unfortunately, no SSD exists whose C achieves these forms because $n<p+1$. Ideally, one would implement a design search algorithm that ranks SSDs according to a summary criterion of the $\Phi_{\lambda}$- or $\Phi_{\lambda}^{\pm}$-criterion. However, these criteria demand intense computations to evaluate a single SSD and search algorithms require many evaluations. We now describe an algorithmic construction that compromises the rigorous but computationally-intensive criteria in Section 2 with new computationally-efficient heuristic criteria based on the theory in Section 3.
[…]
tr(V) and $\text{tr}(\textbf{V})\quad\text{ and }\quad||\textbf{C}-\textbf{C}^{*}||_{F}\ ,$
[…]
To construct designs based on multiple criteria, Cao et al. (2017) combined the coordinate exchange algorithm (Meyer and Nachtsheim, 1995) with an elitist-like operator to produce $ncand$ designs lying near the Pareto front of the criteria. A set of $ntry$ initial designs are provided and each will generate its own Pareto set. That is, for each initial design, many exploratory designs are generated via coordinate exchanges. Only those exploratory designs that either dominate the initial design, or are non-dominated by the other exploratory designs form the initial design’s Pareto set. The procedure continues until each non-dominated design cannot be further improved via coordinate exchange. The approximate Pareto fronts from each initial design are then combined into a single front comprised of the non-dominated designs across all $ntry$ Pareto sets, giving a set of, say, $ncand$ designs. Once the final approximate Pareto set is constructed, we identify which of its $ncand$ designs maximizes the lasso criterion of interest.
The proposed construction algorithm first identifies a set of $ncand$ candidate designs through the Pareto front optimization search algorithm from Cao et al. (2017) based on the two lasso criteria. We then employ the exact lasso sign recovery criteria from Section 2 as a secondary criterion, as it more accurately measures a design’s screening capabilities, to rank the $ncand$ designs and choose the best one. We refer to this algorithm as the Heuristic-Initiated Lasso Sieve, or HILS.
[…]
Figure 6: Plot of the approximate Pareto front for the three designs selected by the coordinate exchange algorithm for the $n=9$, $p=10$, unknown signs case. The starting criteria values for each of the $ntry=20$ starting designs from the algorithm are plotted in black. Each point represents a different design. The criteria values for the DCD are also plotted to show that the approximate Pareto Front designs the DCD in terms of the two criteria.
Although HILS can significantly reduce the number of evaluations of the relaxed lasso criteria compared to constructing an optimal design using, for instance, a coordinate exchange algorithm, there are settings when even a more modest number of evaluations is not computationally feasible. The same issue was encountered by Singh and Stufken (2023). We now discuss some ways to efficiently evaluate the $\Phi_{\lambda}$ and $\Phi_{\lambda}^{\pm}$-criteria, and hence their corresponding summarized versions. For a support size of $k$ and design X, one full evaluation of $\Phi_{\lambda}$ and $\Phi_{\lambda}^{\pm}$ requires consideration of all $\binom{p}{k}$ supports of size $k$. For $\Phi_{\lambda}^{\pm}$, an additional $2^{k-1}$ computations are required for each $\mathcal{A}$ to consider all sign vectors. Evaluating either criteria would benefit from a reduction in the number of supports considered. Sampling from $\mathcal{A}_{k}$ completely at random is intuitive, but may require a large number of samples to be representative. As high correlations between two factors can result in diminished lasso support recovery performance due to low $P(I_{\lambda}\,|\,\textbf{C},\,\boldsymbol{z})$, a support sampling method that balances (or approximately balances) the number of times pairwise sets of factors are included in $\mathcal{A}$ together can be advantageous. Smucker and Drew (2015) utilized nearly balanced incomplete block designs (NBIBDs) to approximate a model space with only a relatively small number of blocks, in the context of model-robust optimal designs. Hence, we recommend implementing the NBIBD sampling method to adequately represent $\mathcal{A}_{k}$ with between 64-128 supports for modest $p$ and $k$ values.
From Lemma 1, for a fixed $\mathcal{A}$, the probabilities are equal between $\boldsymbol{z}_{\mathcal{A}}$ and $-\boldsymbol{z}_{\mathcal{A}}$. Thus, only $2^{k-1}$ sign vectors need be considered for the $\Phi_{\lambda}^{\pm}$-criterion. While this cuts the computation in half, prior knowledge from the practitioner can be leveraged to select an even smaller set of representative sign vectors to evaluate. For example, if there is no knowledge on sign direction, the most likely sign vectors are those with an equal, or nearly equal, number of ±$\pm 1$. For $k=10$, one could reduce the number of sign vectors from $5122^{9}=$ 512 to only the $126$ possible sign vectors having five $+1$’s and $-1$’s. There may also be strong prior belief about the signs of some or all of the factors that can reduce the set of sign vectors to a manageable number, or even to a single element.
5 Examples
We now demonstrate near-optimal SSDs found under the criteria via HILS under three example scenarios. For each case, the coordinate exchange algorithm was executed with $ntry=50$ random starting designs and identified $ncand$ final designs on the approximate Pareto front of the criteria described in Section 4. Of these $ncand$ designs, the final HILS design was selected using the $\Phi_{\text{max}}^{\pm}$ criterion. For each scenario, we provide the computational time that it took to find the HILS-optimal design on a Desktop with an Intel ® Core i7-6700 CPU @ 3.40GHz processor with 16GB RAM. Additionally, we evaluate the performance of the HILS designs compared to competitors under $k$ and $\beta$ settings that are different to the $k$ and $\beta$ values used to construct the HILS designs. This evaluates the sensitivity of the HILS construction method to $k$ and $\beta$.
[…]
5.1 Scenario 1: $n=9$, $p=10$ Unknown Signs
The HILS algorithm was performed for $k=3$ and $\beta=3$ assuming unknown signs. The HILS-optimal design was selected in approximately six minutes of computation time. Because $n\approx p$, this scenario is close to one in which an orthogonal design exists. Therefore, this represents an experiment in which optimal designs under traditional SSD criteria should perform rather well. For unknown signs, we set $c^{*}=0$ following Theorem 3. Since $\textbf{C}^{*}$ is orthogonal and the $\text{UE}(s^{2})$ criterion targets orthogonality directly, the $\text{UE}(s^{2})$-optimal designs were deemed appropriate for comparison. Additionally, since the two criteria used to construct the DCDs are optimized under an orthogonal design, the DCD was also compared.
Figure 7 compares the $\Phi_{\lambda}^{\pm}$ curves as a function of $\log(\lambda)$ for the hypothetically optimal completely symmetric (CS optimal) design, HILS-optimal design, DCD, and a $\text{UE}(s^{2})$-optimal design under a variety of $k$ and $\beta$ settings. The top-middle plot represents the $k$ and $\beta$ setting in which the HILS design was optimized for. The top row of Figure 7, representing $k=3$, demonstrates that the HILS design out performs the others in terms of lasso sign recovery regardless of the $\beta$ value. When $k=5$, Figure 7 shows little difference between the 3 designs. The HILS and $\text{UE}(s^{2})$- optimal designs show a larger $\Phi_{\lambda}^{\pm}$ for larger $\log(\lambda)$, but the DCD attains a slightly larger maximum of $\Phi_{\lambda}^{\pm}$ (for $\beta=3$ and $\beta=4$). Overall, the HILS-optimal design saw better performance than its competitors under the $k$ value that it was optimized for and comparable performance as the $k$ assumption was violated.
Figure 7: $\Phi_{\lambda}^{\pm}$ curves as a function of $\log(\lambda)$ for the HILS optimal design, DCD, and a $\text{UE}(s^{2})$-optimal design under a variety of $k$ and $\beta$ settings for scenario 1. The HILS optimal design shown in all panels was chosen assuming $k=3$ and $\beta=3$. The CS Optimal curve refers to the hypothetical design having the optimal completely symmetric form for $k=3$ and $\beta=3$.
5.2 Scenario 2: $n=9$, $p=10$ Known Signs
To build on the $n$ and $p$ settings in scenario 1, the HILS algorithm was then performed for $k=3$ and $\beta=3$ assuming known signs. Here, the optimal $c$ value was found via numerical optimization to be taken as $c^{*}\approx 0.04$. The HILS algorithm selected the HILS-optimal design in approximately 4 minutes.
Figure 8 compares the $\Phi_{\lambda}$ curves as a function of $\log(\lambda)$ for the CS-optimal design, HILS-optimal design, DCD (Singh and Stufken, 2023), and a $\text{Var}(s+)$-optimal design under a variety of $k$ and $\beta$ settings. To demonstrate the added value these three designs have by incorporating sign information, the figure also has the $\Phi_{\lambda}$ curve for a $\text{UE}(s^{2})$-optimal design. The top-middle plot represents the $k$ and $\beta$ setting in which the HILS design was optimized for. For all $\beta$ and $k$ settings except $\beta=4$ and $k=5$, the HILS-optimal design shows a larger $\Phi_{\text{max}}$ and area under the curve compared to DCD and $\text{Var}(s+)$. For $\beta=4$ and $k=5$, Figure 8 shows a slightly larger $\Phi_{\text{max}}$ for DCD compared to the HILS-optimal design. However, the HILS-optimal design sees higher $\Phi_{\lambda}$ for larger values of $\log(\lambda)$ (around $\log(\lambda)=1$) and has a larger area under its $\Phi_{\lambda}$ curve. Thus, we deem that the HILS-optimal design is superior to the competitors in all settings demonstrated in Figure 8. Finally, we note the poor performance of the $\text{UE}(s^{2})$-optimal design compared to all designs, especially when $k=5$ with $\Phi_{\lambda}\approx 0$. This is most likely due to the prevalence of negative correlations leading to singular $C_{\mathcal{A}}$ for some supports.
Figure 8: $\Phi_{\lambda}$ curves as a function of $\log(\lambda)$ for the HILS-optimal design, DCD, and a $\text{Var}(s+)$-optimal design under a variety of $k$ and $\beta$ settings under scenario 2. The HILS optimal design shown in all panels was chosen assuming $k=3$ and $\beta=3$. The CS Optimal curve refers to the hypothetical design having the optimal completely symmetric form for $k=3$ and $\beta=3$.
5.3 Scenario 3: $n=14$, $p=20$ Known Signs
This scenario assumed known effect directions with $k=5$ and $\beta=2$. The optimal $c$ value was found via numerical optimization to be taken as $c^{*}\approx 0.045$. With a larger difference between $n$ and $p$, the Pareto front found by the coordinate exchange algorithm consisted of 14 designs (the Pareto fronts consisted of 3 designs for Scenarios 1 and 2). The HILS algorithm selected the HILS-optimal design in approximately 1 hour and 21 minutes.
Figure 9 compares the $\Phi_{\lambda}$ curves as a function of $\log(\lambda)$ for the HILS-optimal design, DCD (Singh and Stufken, 2023), and a $\text{Var}(s+)$-optimal design under a variety of $k$ and $\beta$ settings. We again included the $\Phi_{\lambda}$ curve for a $\text{UE}(s^{2})$-optimal design. The top-middle plot represents the $k$ and $\beta$ setting in which the HILS design was optimized for. Across all $\beta$ and $k$ settings there is very little difference between the performances of the three designs that include sign information. However, all of these designs had significantly higher $\Phi_{\lambda}$ values than the $\text{UE}(s^{2})$-optimal design.
Figure 9: $\Phi_{\lambda}$ curves as a function of $\log(\lambda)$ for the HILS-optimal design, DCD, and a $\text{Var}(s+)$-optimal design under a variety of $k$ and $\beta$ settings under scenario 3. The HILS optimal design shown in all panels was chosen assuming $k=5$ and $\beta=2$. The CS Optimal curve refers to the hypothetical design having the optimal completely symmetric form for $k=5$ and $\beta=2$.
The lack of separation between the HILS-optimal design from the other designs was not unexpected given $n$ is nearly half of $p$. It is challenging to find a design that simultaneously minimizes correlations. Nonetheless, the HILS-optimal design performed comparably to the DCD and $\text{Var}(s+)$-optimal designs, with all three significantly outperforming the $\text{UE}(s^{2})$-optimal design. The HILS-optimal design was found much faster than the DCD, which took us dozens of hours to construct. Since the HILS-optimal design can be constructed more quickly and show comparable performance to the DCDs, we view this algorithmic design construction approach as worthwhile.
6 Summary
The SSD literature has predominately constructed designs by optimizing heuristic criteria that measure a design’s proximity to a (nonexistent) orthogonal design. The heuristic criteria are tractable in their optimization, in that optimal designs can generally be constructed directly or algorithmically in a reasonable amount of time. However, these heuristic criteria are not directly tied to a screening analysis method, so there is no guarantee that the resulting analysis under an optimal design will have good statistical properties. This article resolves this disconnect by optimizing criteria based on the probability of sign recovery under the lasso, which is well-defined even when $n-1<p$. The major contributions are:
  1. 1.
    A local optimality criterion assuming known $\boldsymbol{\beta}$ and fixed $\lambda$. A trivial design that confounds all inactive factors with the intercept is shown to be optimal for sign recovery. An exact design construction is given that can improve the probability of sign recovery over an orthogonal design for some $\lambda$ The design has positive and nearly constant pairwise correlations, following the ideal structure for the $\text{Var}(s+)$-criterion.
  2. 2.
    More practical criteria that relax the assumptions about $\boldsymbol{\beta}$ and $\lambda$. Such criteria are computationally intensive and hence difficult to optimize both analytically and algorithmically, requiring computations across all supports of a given size and potentially many sign vectors.
  3. 3.
    A study of the optimal form of C with and without known $\boldsymbol{z}$. By conditioning on completely symmetric matrices, we arrive at a univariate optimization problem. The framework mimics the approximate design approach for least-squares analyses.
  4. 4.
    A rigorous proof that across all completely symmetric matrices, an orthogonal correlation matrix is suboptimal for nearly all $\lambda>0$ when $\boldsymbol{z}$ is known. The optimal form instead takes on positive, constant off-diagonal elements. In the case of unknown sign information, an orthogonal correlation matrix is shown to be a local maximum for a range of $\lambda$. These optimal forms rigorously justify the $\text{UE}(s^{2})$- and $\text{Var}(s+)$-criterion in the cases of unknown and known $\boldsymbol{z}$, respectively. This at least partially solves the two open problems from Section 1.1.
  5. 5.
[…]
The proposed optimal design framework provides an alternative to using simulation to compare SSDs, which can be tedious and difficult to reproduce independently. In particular, there are at least two reasons to be skeptical of conclusions drawn from simulations in the context of SSDs. First, simulations can be misleading when subtly different versions of complicated statistical procedures are used. In our attempts to reproduce and compare the simulation studies of Singh and Stufken (2023) and Weese et al. (2021), we discovered the regularization methods used were sensitive to two different aspects of the procedure implementation. In particular, these papers used the Gauss-Dantzig selector with different ways of exploring the tuning parameter space. They also differed in how they implemented a secondary thresholding of the estimates which determines a subset of the active factors that should remain active. Typically this threshold removes estimates whose absolute magnitude are less than $\sigma^{2}$, but thresholding in this way is of dubious reliability because there is no natural way to estimate $\sigma^{2}$ and thus threshold levels are more or less arbitrary. The approach we present in this paper avoids these difficulties by eliminating the need for simulation at all, in a similar way that closed-form power analysis procedures are routinely used in simple, replicated experimental design settings.
Another danger of the existing heuristic criteria is that there is no guarantee the optimal design has any statistical value. For example, there technically exists a $\text{UE}(s^{2})$-optimal design with $n=2$ runs and $p=100$ factors, but this design has no statistical value. The lasso criteria, however, do reflect statistical value and so can be used to give objective information about design quality, akin to a power analysis. Experimenters can specify a minimum effect size of interest, $\beta$, as well as an educated guess regarding the number of factors that are likely to have such an effect size and investigate the sign-recovery probabilities as a function of $\lambda$. If the maximum average probability is close to 0, it is unlikely that the design can provide reliable sign recovery for the specified effect size. Such a low value could lead to a reconsideration of the runs budget, expected sparsity, and/or the size of effects that are deemed of interest to detect.
The proposed framework allows an even deeper investigation, if desired. For example, we can determine whether a low average probability is due to an inability to reliably exclude all inactive effects (the $I_{\lambda}$ event) or to reliably include all active effects (the $S_{\lambda}$ event). In screening experiments, we are much more concerned with identifying the true effects than with allowing inactive effects through the filter. This suggests that the $S_{\lambda}$ event is more important to the experimenter than the $I_{\lambda}$ event.
There are many avenues of future research to be explored. First, we noted in multiple sections and examples that for larger $\lambda$, $\Phi^{\pm}_{\lambda}(\textbf{C}\,|\,\boldsymbol{\beta})$ sometimes is optimized by a C with large positive off-diagonals, but its improvement is due solely to its ability to recover $\boldsymbol{z}_{\mathcal{A}}=\boldsymbol{1}$. This is practically undesirable because we would like to do well across many possible sign vectors. A similar criticism also holds for averaging across all $\mathcal{A}\in\mathcal{A}_{k}$. That is, the best design may do great for only some supports, and poorly for others. It would be interesting to consider such criteria that consider the minimum probability rather than the average. However, we expect that such criteria would have low probabilities that they would not seem practically useful, as it focuses on the worst possible scenario.
Another area of future work is to explore direct construction techniques for designs with all positive correlations, as such designs may be strong candidates to optimize the known-sign criteria. Continuing with exact design construction, while the HILS algorithm is faster relative to Singh and Stufken (2023), it is much slower than that for the heuristic criteria. More work is needed to make faster design construction algorithms.
Finally, it is important to extend the criteria to incorporate more practical assumptions. Rather than assuming a known magnitude of the active effects, it would be better to incorporate uncertainty of the magnitudes using a prior distribution. This may require even heavier computation, as one would have to integrate the proposed criteria over the prior distribution. We are also considering criteria where the inactive factors have small, but nonzero, effects. This requires a definition of sign recovery with respect to the thresholded lasso in which estimated effects below some threshold value are set to 0.
SUPPLEMENTARY MATERIAL
Following the References are proofs of all results as well as computational details on the evaluation of our criteria. R files that construct designs following (14); replicate the numerical studies in Section 3.1; identify optimal $c$ values for the criteria in Section 3; and construct designs with the HILS algorithm may be found at https://github.com/hkyoung361/Lasso_Optimal_SSD.